### 08 Apr 2026 v1

# Projector, Neural, and Tensor-Network Representations of

Abstract
The $\mathbb{Z}_{N}$ cluster-state wavefunction, a paradigmatic example of symmetry-protected topological (SPT) order with $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ symmetry, is expressed in various equivalent ways. We identify the projector-based scheme called the $P$-representation as the efficient way to express cluster and dipolar cluster state’s wavefunctions. Employing the restricted Boltzmann machine (RBM) scheme to re-write the interaction matrix in the $P$-representation in terms of neural weight matrices allows us to develop the neural quantum state (NQS) and the matrix product state (MPS) representations of the same state. The NQS and MPS representations differ only in the way the weight matrices are split and grouped together in a matrix product. For both $\mathbb{Z}_{N}$ cluster and dipolar cluster states, we derive in closed form the weight function $W(s,h)$ that couples physical spins $s$ to hidden variables $h$, generalizing the previous construction for $\mathbb{Z}_{2}$ cluster states to $\mathbb{Z}_{N}$. For the dipolar cluster state protected by two charge and two dipole symmetries, the procedure we have developed leads to the tensor product state (TPS) representation of the wavefunction where each local tensor carries three virtual indices connecting a given site to two nearest neighbors and one further neighbor. We benchmark the resulting TPS construction against conventional MPS representation using density-matrix renormalization group (DMRG) simulations and argue that the TPS could offer a more efficient representation for some modulated SPT states. As a by-product of the investigation, we generalize the previous $\mathbb{Z}_{2}$ matrix product operator (MPO) construction of the Kramers-Wannier (KW) operator to $\mathbb{Z}_{N}$ and interprets it as the dipolar generalization of the discrete Fourier transform on $\mathbb{Z}_{N}$ variables. The new interpretation naturally explains why the KW map is non-invertible.
I Introduction
Representing many-body quantum states efficiently has been a central theme of condensed matter theory for nearly a century. Starting from the Hartree–Fock approximation, successive developments such as Jastrow-correlated wavefunctions, quantum and variational Monte Carlo, exact diagonalization, and density-matrix renormalization group (DMRG) have all contributed essential insights into strongly correlated states of fermions and bosons. With the advent of modern AI and transformer-based methods, the demand for compact yet accurate descriptions of correlated quantum states has only intensified, together with new opportunities to construct such descriptions [6, 35]. In particular, representing quantum many-body wavefunctions in the spirit of neural networks—most notably via restricted Boltzmann machines (RBM)—has led to the notion of neural quantum states (NQS) [6, 13, 12, 14, 15, 11, 8, 18, 33, 17, 7, 36, 21, 9, 25] which has proven capable of exactly encoding paradigmatic states such as the one-dimensional cluster state and the two-dimensional toric code [13, 12, 14, 8, 9]. In this work, we revisit and systematically generalize the NQS formulation of one-dimensional ${\mathbb{Z}}_{N}$ cluster states, focusing on symmetry-protected topological (SPT) order [10, 29, 26] and its multipolar extensions [16, 20, 19, 34, 24, 28, 4, 23, 2], and clarify its relation to matrix product state (MPS) and more general tensor product state (TPS) constructions. In the process, we develop another efficient way to represent various cluster states based on a local projector, which we name the $P$-representation.
Much progress has been made since the first NQS representation of the ${\mathbb{Z}}_{2}$ cluster state appeared [13, 12, 14, 8]. The ${\mathbb{Z}}_{2}$ cluster state, originally introduced as a universal resource for measurement-based quantum computation [27], has been generalized to ${\mathbb{Z}}_{N}$ cluster states with arbitrary integer $N>2$ and to its variants protected by modulated symmetries—such as dipolar, quadrupolar, and exponential symmetries—which define a broader class of modulated SPT (mSPT) states [16, 20, 28, 22, 37, 4, 2] along with the identification of non-invertible Kramers-Wannier symmetries possessed by the various cluster models [30, 31, 5, 19]. Motivated by these developments, we return to the cluster-state NQS framework with two main goals. First, we extend the NQS construction to uniform ${\mathbb{Z}}_{N}$ and dipole-symmetric ${\mathbb{Z}}_{N}$ cluster models. This is made possible by providing the first closed-form RBM weights for general ${\mathbb{Z}}_{N}$ cluster states. Second, we sharpen the understanding of symmetry fractionalization and and other symmetry aspects of these SPT wavefunctions. Concretely, we construct explicit NQS representations for the ${\mathbb{Z}}_{N}$ cluster state and for two types of dipole-conserving ${\mathbb{Z}}_{N}$ cluster states (type-I and type-II), derive closed-form interaction weight matrices for ${\mathbb{Z}}_{N}$ variables, and show how different ways of splitting the weight matrices yield either an MPS or, in the dipolar case, a TPS with three virtual indices per site. In accomplishing these goals, we introduce the $P$-representation and make extensive use of it as a unifying bridge between NQS and tensor-network-type representations. Using the $P$-reprensentation also helps streamlining proofs of push-through conditions, symmetry fractionalization, and writing down matrix-product-operator (MPO) representation of the ${\mathbb{Z}}_{N}$ Kramers–Wannier (KW) operator. Finally we show both analytically and numerically by performing DMRG calculations that certain dipolar SPT states naturally require TPS as the more efficient way to represent its many-body state. As a by-product of our investigation, we propose the interpretation of the KW operation as the dipolar version of the discrete Fourier transformation and give an intuitive explanation for the non-invertibility of the KW operation.
II $\mathbb{Z}_{N}$ cluster state
We begin with a self-contained review of the ${\mathbb{Z}}_{N}$ cluster state (CS) as the paradigmatic example of SPT phase protected by two global symmetries pertaining to the conservation of ${\mathbb{Z}}_{N}$ charges on even and odd sublattices, respectively. The wavefunction of the CS state represents pairwise interaction between two adjacent ${\mathbb{Z}}_{N}$ spins. Such interacting wavefunction is rendered into what we call the $P$-representation involving the product of local projector $P^{s}$ and the interaction matrix $\bm{\Omega}$. The matrix $\bm{\Omega}$ in turn can be decomposed as a product of two weight matrices, e.g. ${\bm{\Omega}}={\bf W}{\bf W}^{t}$, with the matrix $\bf W$} encoding the interaction between physical (visible) and virtual (hidden) variables. Depending on how the $\bf W$} matrices are split around each local projector, one ends up with different versions of local MPS matrices all leading to the same wavefunction.
II.1 Definition and basic properties
We begin by reviewing basic properties of ${\mathbb{Z}}_{N}$ cluster state. The CS Hamiltonian
$H_{c}=-\sum_{j=1}^{L/2}(Z_{2j-2}^{\dagger}X_{2j-1}Z_{2j}+Z_{2j-1}X_{2j}Z^{\dagger}_{2j+1}+{\rm H.c.})$ (1)
is defined on a closed chain of even length $L$ in terms of generalized Pauli operators $X,Z$ acting on the $\mathbb{Z}_{N}$ basis $|s\rangle$ as
$Z|s\rangle=\omega^{s}|s\rangle,~~X|s\rangle=|s+1\rangle.$ (2)
The model has two $\mathbb{Z}_{N}$ global symmetries [16]:
$C_{1}=\prod_{j}X_{2j-1},~~C_{2}=\prod_{j}X_{2j},$ (3)
representing the sum of ${\mathbb{Z}}_{N}$ charges on the odd and even sublattice, respectively. Its unique ground state $|\psi\rangle=\sum_{\bf s}\Psi(\mathbf{s})|{\bf s}\rangle$, ${\bf s}=\{s_{1},\cdots,s_{L}\}$, has the wavefunction
$\Psi({\bf s})=\omega^{s_{1}s_{2}-s_{2}s_{3}+\cdots-s_{L}s_{1}}$ (4)
with $\omega=\exp(2\pi i/N)$. The invariance of $\Psi({\bf s})$ under the $C_{1}$ symmetry operation $s_{2j-1}\rightarrow s_{2j-1}-1$ or the $C_{2}$ operation $s_{2j}\rightarrow s_{2j}-1$ is readily checked.
The same wavefunction $\Psi({\bf s})$ can be cast as MPS:
$\Psi({\bf s})={\rm Tr}[A^{s_{1}}\cdots A^{s_{L}}],$ (5)
provided the local tensors are chosen as
$A^{s_{2j-1}}$ $=\sum_{g\in{\mathbb{Z}}_{N}}\omega^{-s_{2j-1}g}|g\rangle\langle s_{2j-1}|\,$,
$A^{s_{2j}}$ $=\sum_{g\in{\mathbb{Z}}_{N}}\omega^{s_{2j}g}|g\rangle\langle s_{2j}|\,.$ (6)
Invariance of $\Psi({\bf s})$ under $C_{1}^{\dagger},C_{2}^{\dagger}$ can be checked in its MPS representation by observing that changing $s\rightarrow s+1$ results in
$A^{s_{2j-1}+1}$ $=Z^{\dagger}A^{s_{2j-1}}X^{\dagger},$
$A^{s_{2j}+1}$ $=ZA^{s_{2j}}X^{\dagger},$ (7)
in a manifestation of the push-through condition of MPS. The wavefunction itself undergoes the change
$\Psi({\bf s})$ $\xrightarrow{C_{1}^{\dagger}}{\rm Tr}[\cdots A^{s_{2j-1}+1}A^{s_{2j}}A^{s_{2j+1}+1}\cdots]$
$={\rm Tr}[\cdots(Z^{\dagger}A^{s_{2j-1}}X^{\dagger})A^{s_{2j}}(Z^{\dagger}A^{s_{2j+1}}X^{\dagger})\cdots],$ (8)
where we used (II.1) in the second line. The various virtual $X$ operations in the second line can be re-grouped as
${\rm Tr}[\cdots Z^{\dagger})A^{s_{2j-1}}(X^{\dagger}A^{s_{2j}}Z^{\dagger})A^{s_{2j+1}}(X^{\dagger}A^{s_{2j+2}}Z^{\dagger})\cdots].$
The identity $X^{\dagger}A^{s_{2j}}Z^{\dagger}=A^{s_{2j}}$ obeyed by even-site MPS matrices ensures that this is equal to the original CS wavefunction. Another way to see the restoration of the wavefunction is to note that, taking a pair of matrices $A^{s_{2j}}A^{s_{2j+1}}$ as a unit, the transformation under $C_{1}^{\dagger}$ gives
$A^{s_{2j}}A^{s_{2j+1}}\xrightarrow{C_{1}^{\dagger}}A^{s_{2j}}Z^{\dagger}A^{s_{2j+1}}X^{\dagger}=X(A^{s_{2j}}A^{s_{2j+1}})X^{\dagger}.$ (9)
The auxiliary matrices $X$ and $X^{\dagger}$ cancel each other out in taking the product over all sites. A similar exercise using the identity $X^{\dagger}A^{s_{2j-1}}Z=A^{s_{2j-1}}$ for the odd-site matrices shows that the wavefunction is preserved under $C_{2}^{\dagger}$ as well.
Next we turn to the open chain of even length $L$, and consider the CS Hamiltonian:
$H=$ $-Z_{2}^{\dagger}X_{3}Z_{4}-Z_{1}X_{2}Z^{\dagger}_{3}$-$\cdots$
$-Z_{L-2}^{\dagger}X_{L-1}Z_{L}-Z_{L-3}X_{L-2}Z^{\dagger}_{L-1}+{\rm H.c.}.$ (10)
Ground states of this model are
$|\Psi(s_{1},s_{L})\rangle=\sum_{{\bf s}^{\prime}}\langle s_{1}|A^{s_{2}}\cdots A^{s_{L}}|s_{L}\rangle|{\bf s}\rangle,$ (11)
with the ${\bf s}^{\prime}=\{s_{2},\cdots s_{L-1}\}$ and fixed $\{s_{1},s_{L}\}$. For this state we show in App. A
$C_{1}^{\dagger}|\Psi(s_{1},s_{L})\rangle$ $=X_{1}Z_{2}Z_{L}^{\dagger}|\Psi(s_{1},s_{L})\rangle\,$,
$C_{2}^{\dagger}|\Psi(s_{1},s_{L})\rangle$ $=Z_{1}^{\dagger}Z_{L-1}X_{L}|\Psi(s_{1},s_{L})\rangle\,$, (12)
demonstrating how the global symmetries $C_{1}^{\dagger}$ and $C_{2}^{\dagger}$ fractionalize into edge operators $L_{i},R_{i},\,i=1,2$:
$C_{1}^{\dagger}$ $\mapsto L_{1}R_{1},\quad(L_{1},R_{1})=(X_{1}Z_{2},Z_{L}^{\dagger}),$
$C_{2}^{\dagger}$ $\mapsto L_{2}R_{2},\quad(L_{2},R_{2})=(Z_{1}^{\dagger},Z_{L-1}X_{L}).$ (13)
While $C_{1}^{\dagger},C_{2}^{\dagger}$ commute, their fractionalizations commute only projectively:
$L_{1}L_{2}=\omega L_{2}L_{1},\quad R_{2}R_{1}=\omega R_{1}R_{2},$ (14)
generating the NN-fold degeneracy at each edge. The corresponding eigenstates can be constructed explicitly as, e.g. $|\Psi(s_{1},s_{L})\rangle=L_{1}^{s_{1}}R_{2}^{s_{L}}|\Psi(0,0)\rangle$, using $\{s_{1},s_{L}\}$ as quantum numbers labeling each ground state.
II.2 CS wavefunction in $P$, NQS, and MPS representations
We begin with a simple observation that the CS wavefunction becomes
$\Psi({\bf s})={\rm Tr}[P^{s_{1}}{\bm{\Omega}}P^{s_{2}}{\bm{\Omega}}^{*}P^{s_{3}}{\bm{\Omega}}\cdots P^{s_{L}}{\bm{\Omega}}^{*}],$ (15)
where $P^{s}=|s\rangle\langle s|$ is the projector and the interaction matrix $\bm{\Omega}$ is
${\bm{\Omega}}$ $=\sum_{g_{1},g_{2}\in{\mathbb{Z}}_{N}}\omega^{g_{1}g_{2}}|g_{1}\rangle\langle g_{2}|.$ (16)
Note that $g_{1},g_{2}$ are virtual, or bond ${\mathbb{Z}}_{N}$ variables. Since the physical spins appear only through the projector $P^{s}$ in Eq. (15), we refer to it as the $P$-representation.
In turn, the interaction $\omega^{ss^{\prime}}$ between adjacent spins can be thought of as the effective interaction emerging from the interaction of physical spins $s,s^{\prime}$ with a common hidden variable $h$. Explicitly, one can look for the RBM reprensetation
$\sum_{h\in{\mathbb{Z}}_{N}}W(s_{1},h)W(s_{2},h)=\omega^{s_{1}s_{2}},$ (17)
fulfilled by some weight function $W(s,h)$. The explicit construction of $W(s,h)$ for ${\mathbb{Z}}_{2}$ CS was found before [13, 14]. Here we provide its ${\mathbb{Z}}_{N}$ generalization in closed form:
$W(s,h)$ $=\kappa^{-1/2}\omega^{ah^{2}+bs^{2}+csh}$
$\kappa$ $=\sum_{h=0}^{N-1}\omega^{2ah^{2}},$ (18)
with the coefficients
$a=\frac{N-1}{4}\,,\quad b=\frac{N-1}{2}\,,\quad c=-1\,.$ (19)
For even NN, it suffices to choose (a,b,c)=(1/4,1/2,1)(a,b,c)=(-1/4,-1/2,-1). For derivation, see App. B. We can write (II.2) as a matrix relation:
${\bf W}$ $=\sum_{g,h\in{\mathbb{Z}}_{N}}W(g,h)|g\rangle\langle h|$
${\bf W}^{t}$ $=\sum_{g,h\in{\mathbb{Z}}_{N}}W(g,h)|h\rangle\langle g|$
${\bm{\Omega}}$ $={\bf W}{\bf W}^{t}.$ (20)
Plugging the last line ${\bm{\Omega}}={\bf W}{\bf W}^{t}$ into the $P$-representation (15) results in the NQS representation of the CS wavefunction:
$\Psi({\bf s})={\rm Tr}[P^{s_{1}}{\bf W}{\bf W}^{t}P^{s_{2}}{\bf W}^{*}{\bf W}^{\dagger}P^{s_{3}}{\bm{\Omega}}\cdots P^{s_{L}}{\bf W}^{*}{\bf W}^{\dagger}].$ (21)
Finally, we arrive at the MPS representation of the CS wavefunction by noting that Eq. (15) can be written as a matrix product ${\rm Tr}[A^{s_{1}}A^{s_{2}}\cdots]$ provided we set
$A^{s_{2j-1}}={\bm{\Omega}}^{*}P^{s_{2j-1}},~~A^{s_{2j}}$ $={\bm{\Omega}}P^{s_{2j}}.$ (22)
This is precisely the MPS matrix given earlier in Eq. (6). There is an alternative way to derive the matrix $A^{s}$ which starts from the NQS representation in (21) and groups terms as
$A^{s_{2j-1}}$ $={\bf W}^{\dagger}P^{s_{2j-1}}{\bf W}\,$,
$A^{s_{2j}}$ $={\bf W}^{t}P^{s_{2j}}{\bf W}^{*}\,.$ (23)
The matrix elements of these MPS are
$A^{s_{2j-1}}_{\alpha\beta}$ $=\frac{1}{|\kappa|}\omega^{\frac{1}{4}(\alpha^{2}-\beta^{2})+s_{2j-1}(\alpha-\beta)}\,$,
$A^{s_{2j}}_{\alpha\beta}$ $=\frac{1}{|\kappa|}\omega^{-\frac{1}{4}(\alpha^{2}-\beta^{2})-s_{2j-1}(\alpha-\beta)}\,.$ (24)
We began by writing the CS wavefunction in the $P$-representation as the (trace of) the product of projectors and interaction matrices, then performed decomposition of the interaction matrix as a product of two weight matrices to arrive at its NQS representation. The MPS representation of the CS wavefunction follows by identifying the local MPS matrix with the product of a projector and an interaction matrix. Depending on how the identification goes, we may have more than one way to write the MPS matrix, e.g. Eq. (22) and Eq. (23), that are unitarily equivalent. The MPS=NQS correspondence demonstrated here for the ${\mathbb{Z}}_{N}$ cluster state is consistent with earlier observations relating NQS and tensor-network states [15, 11].
II.3 Symmetry transformations
It has been a common practice to prove the invariance of the CS wavefunction under the global symmetries $C_{1},C_{2}$, as well as their fractionalization in an open chain, in the framework of MPS representation. Here we show that the same analysis can be performed just as easily, if not more so, using the $P$-representation.
The symmetry transformation of the projector $P^{s}\rightarrow P^{s+1}$ results in
$P^{s+1}=XP^{s}X^{\dagger},$ (25)
which is in fact the push-through condition. Feeding this change into (15) gives
$\Psi(\mathbf{s})$ $\xrightarrow{C_{1}^{\dagger}}\textrm{Tr}\left[(XP^{s_{1}}X^{\dagger})\bm{\Omega}P^{s_{2}}\bm{\Omega}^{*}(XP^{s_{3}}X^{\dagger})\cdots P^{s_{L}}\bm{\Omega}^{*}\right]$
$=\textrm{Tr}\left[P^{s_{1}}(X^{\dagger}\bm{\Omega}P^{s_{2}}\bm{\Omega}^{*}X)P^{s_{3}}\cdots(X^{\dagger}\bm{\Omega}P^{s_{L}}\bm{\Omega}^{*}X)\right]$ (26)
after some re-grouping of terms. The matrix $\Omega$ has some easily verified properties
$X^{\dagger}\bm{\Omega}$ $=\bm{\Omega}Z\,,~~\bm{\Omega}^{*}X=Z^{\dagger}\bm{\Omega}^{*}\,$,
$\bm{\Omega}X$ $=Z\bm{\Omega}\,,~~X^{\dagger}\bm{\Omega}^{*}=\bm{\Omega}^{*}Z^{\dagger}\,$, (27)
that can be used to convert the expression in the second line of (26) to
$\textrm{Tr}\left[P^{s_{1}}\bm{\Omega}(ZP^{s_{2}}Z^{\dagger})\bm{\Omega}^{*}P^{s_{3}}\cdots\bm{\Omega}(ZP^{s_{L}}Z^{\dagger})\bm{\Omega}^{*}\right].$ (28)
Each term inside the parenthesis obeys $ZP^{s}Z^{\dagger}=P^{s}$, establishing the equivalence of the transformed wavefunction to the original one. The symmetry of the wavefunction under $C_{2}$ can be proven similarly. The algebraic proof given here can be equivalently illustrated as a graphical proof - see Fig. 1.
Figure 1: Graphical proof of the invariance of the CS wavefunction under $C_{1}$ in the $P$-representation. The symmetry operation by $X$ on the projector $P^{s}$ at the odd sites shown in (a) becomes virtual operations on the same projector as shown in (b). The left (right) action on $\bm{\Omega}$ ($\bm{\Omega}^{*}$) becomes the right (left) action on $\bm{\Omega}$ ($\bm{\Omega}^{*}$) as shown in (c). The net result is the conjugation of the even-site projector by $Z$ and $Z^{\dagger}$ as shown in (d), which gives back the original projector and recovers the CS wavefunction.
For the open chain of length $L$, the $P$-representation of the CS wavefunction becomes
$|\Psi(s_{1},s_{L})\rangle=\sum_{{\bf s}^{\prime}}\langle s_{1}|{\bm{\Omega}}P^{s_{2}}{\bm{\Omega}}^{*}P^{s_{3}}{\bm{\Omega}}\cdots P^{s_{L-1}}\bm{\Omega}|s_{L}\rangle|\mathbf{s}\rangle,$ (29)
with a pair of fixed edge spins $\{s_{1},s_{L}\}$. This is identical to the MPS representation on an open chain given in Eq. (11). The proof of symmetry fractionalization, shown in Eq. (12), automatically follows.
The invariance of the CS state under global symmetries can be checked explicitly in the RBM-inspired MPS representation, Eq. (23). The proof, reproduced in App. C, is quite a bit more complicated than what is required to prove the same statement in the $P$-representation.
III Dipolar cluster state
Dipolar and other multipolar cluster states have been constructed as part of an effort to construct SPT states protected by spatially modulated symmetries. Here we discuss two types of dipolar CS as representative examples of mSPT and discuss their $P$, NQS, and MPS representations.
III.1 Type-I dipolar cluster state
The dipolar cluster state (dCS) was introduced [16] as an example of SPT protected by one uniform and one dipole-modulated symmetries
$C=\prod_{j}X_{j},~~D=\prod_{j}(X_{j})^{j}.$ (30)
An explicit model Hamiltonian can be constructed with these symmetries,
$H=-\sum_{j=1}^{L}(Z_{j-1}Z^{\dagger}_{j}X_{j}Z^{\dagger}_{j}Z_{j+1}+h.c.),$ (31)
where $L$ must be divisible by NN for a closed chain to unambiguously define the dipole symmetry operator $D$. Its unique ground state on a periodic chain is [16]
$\Psi({\bf s})$ $=\omega^{\sum_{j=1}^{L}s_{j}(s_{j+1}-s_{j})}={\rm Tr}[A^{s_{1}}\cdots A^{s_{L}}],$
$A^{s}$ $=\omega^{-s^{2}}\sum_{g}\omega^{sg}|g\rangle\langle s|.$ (32)
The invariance of this wavefunction under $C$ and $D$ can be readily confirmed. In particular, the $C$-invariance of the MPS wavefunction follows directly from the relation
$A^{s+1}=(ZX)A^{s}(ZX)^{\dagger}$
for the matrix $A^{s}$ in (32), which can be verified by an explicit calculation. Under the $D$ transformation,
${\rm Tr}[A^{s_{1}}\cdots A^{s_{L}}]$ $\xrightarrow{D}{\rm Tr}[A^{s_{1}}ZXA^{s_{2}}ZX\cdots A^{s_{L}}(X^{\dagger}Z^{\dagger})^{L}ZX]$
$=(-1)^{L+L/N}{\rm Tr}[(A^{s_{1}}ZX)(A^{s_{2}}ZX)\cdots(A^{s_{L}}ZX)]$
where we invoked the identity $(X^{\dagger}Z^{\dagger})^{N}=(-1)^{N+1}$ in the second line. Further using the identity $ZXA^{s}=ZA^{s}Z^{\dagger}$ restores the original wavefunction, up to a phase factor. This line of proof of invariance was first presented in [16] and is reproduced here for completeness. This dipolar cluster state is referred to as type-I, or dCSI, to distinguish it from the type-II dCS state to be defined shortly.
The dCSI wavefunction in the $P$-representation is simply
$\Psi({\bf s})$ $={\rm Tr}[P^{s_{1}}{\bm{\Omega}}P^{s_{2}}{\bm{\Omega}}\cdots P^{s_{L}}{\bm{\Omega}}],$
$P^{s}$ $=\omega^{-s^{2}}|s\rangle\langle s|,$ (33)
using the same interaction matrix $\bm{\Omega}$ previously introduced in (20), and the modified projector $P^{s}$ containing an extra factor $\omega^{-s^{2}}$. The NQS representation follows from inserting ${\bm{\Omega}}={\bf W}{\bf W}^{t}$ in the $P$-representation. The MPS representation, in turn, follows from $A^{s}=P^{s}{\bm{\Omega}}$, which reproduces (32). Alternatively, one can write the MPS matrix as $A^{s}={\bf W}^{t}P^{s}{\bf W}$. Explicitly,
$A^{s}_{\alpha\beta}=\frac{1}{\kappa}\omega^{-\frac{1}{4}(\alpha^{2}+\beta^{2})-2s^{2}-s(\alpha+\beta)}\,.$ (34)
The proof of invariance of the dCSI wavefunction under $C$ and $D$ proceeds by noting that the modified projector $P^{s}$ obeys the push-through condition
$P^{s+1}=Z^{\dagger}XP^{s}Z^{\dagger}X^{\dagger}=XZ^{\dagger}P^{s}X^{\dagger}Z^{\dagger}.$ (35)
The dCSI wavefunction accordingly transforms as
$\Psi(\mathbf{s})\xrightarrow{C^{\dagger}}$ ${\rm Tr}[P^{s_{1}}(X^{\dagger}Z^{\dagger}\bm{\Omega}XZ^{\dagger})P^{s_{2}}(X^{\dagger}Z^{\dagger}\bm{\Omega}XZ^{\dagger})\cdots$
$\quad\cdots P^{s_{L}}(X^{\dagger}Z^{\dagger}\bm{\Omega}XZ^{\dagger})],$ (36)
where we introduced parentheses to group terms around each $\bm{\Omega}$. Since $\bm{\Omega}$ satisfies its own push-through condition
$X^{\dagger}Z^{\dagger}\bm{\Omega}XZ^{\dagger}=\bm{\Omega},$ (37)
the invariance of the wavefunction $\Psi({\bf s})$ under $C$ is established. The same proof can be performed step by step in a graphical manner, as illustrated in Fig. 2.
Figure 2: Graphical proof of the invariance of the dCSI wavefunction under $C$ in the $P$-representation. Symmetry operation by $X$ on the physical index of each projector $P^{s_{j}}$ shown in a (a) leads to operations on their virtual indices as shown in (b). Viewed as operations on $\bm{\Omega}$ from the left and the right, they restore the original wavefunction by the push-through condition (37).
Figure 3: Graphical proof of the invariance of the dCSI wavefunction under $D$ in the $P$-representation. Symmetry operation by $X^{j}$ on the physical index of each projector $P^{s_{j}}$ shown in (a) leads to operations on their virtual indices as shown in (b). The resulting structure can be viewed as operations on $\bm{\Omega}$ from the left and the right. Exploiting the push-through condition (37) repeatedly, most of the terms cancel out except for one factor of $XZ^{\dagger}$ on the right side of each $\bm{\Omega}$ as shown in (c). Finally, the $X$ on the right side of $\bm{\Omega}$ transforms to $Z$ on its left according to (27). The only remaining term is $Z$ conjugating each projector $P^{s_{j}}$ as shown in (d), which restores the original wavefunction since $Z^{\dagger}P^{s}Z=P^{s}$.
The proof of invariance under $D$ proceeds by first noting the push-through condition
$P^{s_{j}+j}=(Z^{\dagger}X)^{j}P^{s_{j}}(Z^{\dagger}X^{\dagger})^{j}=(XZ^{\dagger})^{j}P^{s_{j}}(X^{\dagger}Z^{\dagger})^{j}.$ (38)
Rather than going through the rest of the derivation algebraically, we refer to the graphical proof in Fig. 3. That argument reproduces the result $\Psi(\mathbf{s})\xrightarrow{D^{\dagger}}(-1)^{L+L/N}\,\Psi(\mathbf{s})$ from the MPS representation up to the overall sign; The correct sign is then recovered by taking into account $(X^{\dagger}Z^{\dagger})^{L}=(-1)^{L+L/N}$. Other aspects of the dCSI such as the symmetry fractionaliztion can be worked out by going through similar analyses as before.
III.2 Type-II dipolar cluster state: $P$-, NQS, and TPS representations
A second type of dipolar cluster model was proposed in [19] with the Hamiltonian
$H=-\sum_{j}Z_{2j-1}X_{2j}Z_{2j+1}^{-2}Z_{2j+3}-\sum_{j}Z_{2j-2}Z_{2j}^{-2}X_{2j+1}Z_{2j+2}+h.c.,$ (39)
which is is symmetric under two sets of charge and dipole symmetries
$C_{1}$ $=\prod_{j}X_{2j-1},~~C_{2}=\prod_{j}X_{2j}$
$D_{1}$ $=\prod_{j}(X_{2j-1})^{j},~~D_{2}=\prod_{j}(X_{2j})^{j}.$ (40)
To distinguish this from the previous dipolar CS, we refer to the new model as dCSII. The ground-state wavefunction of the dCSII Hamiltonian on a closed chain is [19]:
$\Psi({\bf s})$ $=\prod_{j}\omega^{s_{2j}(s_{2j-1}-2s_{2j+1}+s_{2j+3})}$
$=\prod_{j}\Omega_{s_{2j-1},s_{2j}}\tilde{\Omega}_{s_{2j},s_{2j+1}}\Omega_{s_{2j},s_{2j+3}}\,$, (41)
where $\Omega_{ss^{\prime}}=\omega^{ss^{\prime}}$ and $\tilde{\Omega}_{ss^{\prime}}=\omega^{-2ss^{\prime}}$.
To cast the wavefunction in the $P$-representation, we introduce the projector $P^{s}$ with three virtual indices,
$P^{s}_{\alpha\beta\gamma}=\delta_{s\alpha}\delta_{s\beta}\delta_{s\gamma}.$ (42)
The dCSII wavefunction then becomes
$\Psi(\mathbf{s})$ $=\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\tilde{\Omega}_{\beta^{\prime}_{j}\beta_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}\,$, (43)
with ${\bm{\alpha}}=\{\alpha_{1},\cdots\alpha_{L}\}$, etc.. It is illustrated graphically in Fig. 4 (a). By appropriately organizing terms around each site, we can directly obtain a TPS representation with the local tensors carrying three virtual indices:
$\Psi(\mathbf{s})$ $=\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\prod_{j}T^{s_{2j}}_{\alpha_{j}\beta_{j}\gamma_{j}}T^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}\,$,
$T^{s_{2j}}_{\alpha_{j}\beta_{j}\gamma_{j}}$ $=\sum_{\alpha^{\prime}_{j}\gamma^{\prime}_{j}}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}\,$,
$T^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$ $=\sum_{\beta^{\prime}_{j}}P^{s_{2j+1}}_{\alpha_{j+1}\beta^{\prime}_{j}\gamma_{j-1}}\tilde{\Omega}_{\beta_{j}\beta^{\prime}_{j}}\,.$ (44)
To cast the wavefunction in the NQS representation, we introduce a second weight function $\tilde{W}(s,h)$ satisfying
$\sum_{h}\tilde{W}(s_{1},h)\tilde{W}(s_{2},h)$ $=\omega^{-2s_{1}s_{2}}\,$, (45)
in addition to $W(s,h)$ given in (17). Such function can be found
$\tilde{W}(s,h)$ $=\tilde{\kappa}^{-1/2}\omega^{\tilde{a}h^{2}+\tilde{b}s^{2}+\tilde{c}sh}$
$\tilde{\kappa}$ $=\sum_{h=0}^{N-1}\omega^{2\tilde{a}h^{2}}$ (46)
with
$(\tilde{a},\tilde{b},\tilde{c})=\left(\frac{1}{2},1,2\right)\,$, (47)
for $N\neq 2\mod 4$. For $N=2\mod 4$, the constants are modified to
$(\tilde{a},\tilde{b},\tilde{c})=\left(\frac{N+2}{4},1+\frac{N}{2},2\right)\,.$ (48)
For derivation of the coefficients, see App. B. Plugging the $W$’s and $\tilde{W}$’s into (III.2) gives the NQS representation of the dCSII wavefunction. It is shown graphically in Fig. 4 (b).
One can group the various $W$ and $\tilde{W}$ weights around each site and arrive at an alternative TPS representation:
$\Psi(\mathbf{s})=\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\prod_{j}T^{s_{2j}}_{\alpha_{j}\beta_{j}\gamma_{j}}T^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}},$ (49)
where
$T^{s_{2j}}_{\alpha_{j}\beta_{j}\gamma_{j}}$ $=W(s_{2j},\alpha_{j})\tilde{W}(s_{2j},\beta_{j})W(s_{2j},\gamma_{j})$
$=\frac{1}{\sqrt{\kappa^{2}\tilde{\kappa}}}\omega^{-\frac{1}{4}(\alpha_{j}^{2}+\gamma_{j}^{2}-2\beta_{j}^{2})-s_{2j}(\alpha_{j}+\gamma_{j}-2\beta_{j})}$
$T^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$ $=W(s_{2j+1},\alpha_{j+1})\tilde{W}(s_{2j+1},\beta_{j})W(s_{2j+1},\gamma_{j-1})$
$=\frac{1}{\sqrt{\kappa^{2}\tilde{\kappa}}}\omega^{-\frac{1}{4}(\alpha_{j+1}^{2}+\gamma_{j-1}^{2}-2\beta_{j}^{2})-s_{2j+1}(\alpha_{j+1}+\gamma_{j-1}-2\beta_{j})}.$ (50)
Figure 4: (a) $P$-representation of the dCSII wavefunction. (b) NQS representation of the dCSII wavefunction. Blue and pink circles indicate visible variables on the even and odd sites respectively. The yellow and purple dots represent connections between hidden and visible units by $W$ and $\tilde{W}$ respectively. (c) TPS representation of the dCSII wavefunction. Blue and pink squares indicate the tensors $T^{s_{2j}}$ and $T^{s_{2j+1}}$ respectively. Each tensor has one physical and three virtual bonds.
Let us now show that the dCSII wavefunction is invariant under all four symmetry operations $C_{1},C_{2}$ and $D_{1},D_{2}$, using its $P$-representation in (43). As before, such proof begins by expressing how the physical action $s\rightarrow s+1$ in the projector $P^{s}_{\alpha\beta\gamma}$ is transferred to the virtual degrees of freedom. The proof is most easily demonstrated graphically, as illustrated in Figs. 5 and 6. The algebraic proof, equivalent to the graphical one, can be found in App. D.
A proof of symmetry based on the NQS representation of the dCSII wavefunction is possible, by leveraging some transformation properties of the weight functions ($k\in{\mathbb{Z}}_{N}$):
$W(s+k,h)$ $=W(s,h)\omega^{2bks+ckh+bk^{2}}\,$,
$\sum_{h}W(x,h)W(y,h)\omega^{ckh}$ $=\omega^{-bk^{2}}\omega^{-2bxy}\omega^{-2bk(x+y)}\,.$ (51)
The $W$ and $\tilde{W}$ weights used in the NQS construction of the dCSII both belong to this family of functions with specific choices of constants $(a,b,c)$. As detailed in App. D, these relations are sufficient to prove
$\Psi(\mathbf{s})\xrightarrow[]{C_{1},C_{2},D_{1},D_{2}}\Psi(\mathbf{s})\,.$ (52)
Figure 5: Graphical proof of the invariance of the dCSII wavefunction under $C_{2}$ in the $P$-representation. The symmetry operation by $X$ on the projector $P^{s}$ at the even sites shown in (a) becomes virtual operations on the same projector as shown in (b). The virtual $X$ operations are moved through the interaction matrices $\bm{\Omega}$, $\tilde{\bm{\Omega}}$, which convert them into $Z$ operations as shown in (c). The resultant operations are $Z$ operations around the projectors at odd sites whose sum of exponents is zero as shown in (d), thereby giving the original wavefunction.
Figure 6: Graphical proof of the invariance of the dCSII wavefunction under $D_{2}$ in the $P$-representation. The symmetry operation by $X^{j}$ on the projector $P^{s}$ at the even sites shown in (a) becomes virtual operations on the same projector as shown in (b). The virtual $X^{j}$ operations are moved through the interaction matrices $\bm{\Omega}$, $\tilde{\bm{\Omega}}$, which convert them into $Z$ operations as shown in (c). The resultant operations are $Z$ operations around the projectors at odd sites whose sum of exponents is zero as shown in (d), thereby giving the original wavefunction.
A general lesson may be drawn from our endeavor. Modulated SPTs as exemplified by the dCSII state has wavefunctions that reflect interactions among spins going beyond the nearest neighbor. As long as the interactions are of pairwise nature, though, the wavefunction can be readily cast in the $P$-representation, which in turn facilitates the analysis of symmetry invariance and its fractionalization through push-through conditions of the $P$-tensor. Furthermore, the $P$-representation generalizes easily to projectors having multiple virtual legs, resulting in the TPS representation of the wavefunction. Other mSPTs such as those protected by quadrupolar and exponential symmetries [16, 23, 19] are amenable to similar $P$-reprensetation analysis, though we do not pursue the issues here.
III.3 Generality of the $P$-representation
Since we placed much emphasis on the $P$-representation, it is worth going over its utility in representing more general quantum states. Consider a many-body wavefunction with ${\mathbb{Z}}_{N}$ spins ${\bf s}=\{s_{1},\cdots,s_{L}\}$ given by arbitrary pairwise interaction
$\Psi({\bf s})=\Omega_{12}(s_{1},s_{2})\Omega_{23}(s_{2},s_{3})\cdots,$ (53)
with the subscript to emphasize that each pairwise function $\Omega_{j,j+1}(s_{j},s_{j+1})$ may differ from pair to pair. All such wavefunctions can be written in the $P$-reprensetation:
$\Psi({\bf s})={\rm Tr}[P^{s_{1}}{\bm{\Omega}}_{12}P^{s_{2}}{\bm{\Omega}}_{23}\cdots]$ (54)
with the interaction matrix $\bm{\Omega}$
${\bm{\Omega}}_{j,j+1}=\sum_{g_{j},g_{j+1}}\Omega_{j,j+1}(g_{j},g_{j+1})|g_{j}\rangle\langle g_{j+1}|.$ (55)
The $g_{j},g_{j+1}$ refer to virtual, or bond degrees of freedom. One can interpret the $P$-representation in (54) as MPS by associating
$A^{s_{j}}=P^{s_{j}}{\bm{\Omega}}_{j,j+1}~~{\rm or}~~A^{s_{j}}={\bm{\Omega}}_{j-1,j}P^{s_{j}}.$ (56)
The NQS representation follows from decomposing the interaction matrix $\bm{\Omega}$}
$\Omega_{j,j+1}(g_{j},g_{j+1})=\sum_{h_{j,j+1}\in{\mathbb{Z}}_{N}}L_{j}(g_{j},h_{j,j+1})R_{j+1}(h_{j,j+1},g_{j+1})$ (57)
for some weight functions $L$ and $R$. More compactly,
${\bm{\Omega}}_{j,j+1}$ $={\bf L}_{j}{\bf R}_{j+1}$
${\bf L}_{j}$ $=\sum_{g_{j},h_{j,j+1}}L_{j}(g_{j},h_{j,j+1})|g_{j}\rangle\langle h_{j,j+1}|$
${\bf R}_{j+1}$ $=\sum_{h_{j,j+1},g_{j+1}}R_{j+1}(h_{j,j+1},g_{j+1})|h_{j,j+1}\rangle\langle g_{j+1}|.$ (58)
Inserting the decomposition ${\bm{\Omega}}_{j,j+1}={\bf L}_{j}{\bf R}_{j+1}$ into the $P$-representation in (54) gives
${\rm Tr}[P^{s_{1}}{\bf L}_{1}{\bf R}_{2}P^{s_{2}}{\bf L}_{2}{\bf R}_{3}\cdots]$ (59)
which gives rise to the the NQS-inspired MPS matrix
$A_{j}^{s_{j}}={\bf R}_{j}P^{s_{j}}{\bf L}_{j}.$ (60)
The expression of MPS matrix $A^{s}$ as the conjugation of the projector $P^{s}$ with a pair of matrices ${\bf L},{\bf R}$ is a general feature of wavefunctions given by (53).
Writing the interaction matrix ${\bm{\Omega}}_{j,j+1}$ as the product ${\bf L}_{j}{\bf R}_{j+1}$ is a singular value decomposition (SVD) problem. For a given matrix $\bm{\Omega}$} there is a corresponding SVD ${\bm{\Omega}}={\bf U}{\bm{\Sigma}}{\bf V}^{*}$ and by associating ${\bf L}={\bf U}{\bm{\Sigma}}^{1/2}$ and ${\bf R}={\bm{\Sigma}}^{1/2}{\bf V}^{*}$, one arrives at the desired factorization. A trivial decomposition with ${\bf L}=\mathbb{I}$ or ${\bf R}=\mathbb{I}$ with the other matrix being equal to $\bm{\Omega}$} results in the corresponding MPS given by either $A^{s}=P^{s}{\bm{\Omega}}$ or $A^{s}={\bm{\Omega}}P^{s}$.
As we witnessed in the case of dCSII wavefunction, the utility of $P$-representation extends beyond the case of nearest-neighbor weights assumed in (53). In that case, however, the projector must be generalized to have virtual indices equal to the number of neighbors to which a given spin $s$ is connected.
III.4 Representational aspects of NQS and TPS
So far our work focused on demonstrating the equivalence of MPS to the NQS representation of the same state embodying various symmetry-protected orders. This has been particularly transparent in the case of cluster state and the type-I cluster state, where the NQS representations gives a nice behind-the-scenes interpretation of where the MPS representation with its hidden units come from. In short, the NQS leads to MPS by re-organizing the weight functions around a site rather than around a bond.
The MPS=NQS connection becomes more intricate for the type-II dCS state, where we do not really have an MPS representation of the state in the proper sense of the word, but the NQS representation can be obtained more straightforwardly. Specifically, the weight functions $W$ and $\tilde{W}$ can be derived directly from the pairwise structure of the wavefunction in Eq. (III.2), and the re-organization of these weight functions around each site naturally gives rise to the TPS with rank-3 tensors in Eq. (III.2). Arriving at such a tensor network ansatz from the bare wavefunction alone would be considerably less obvious. We believe this conceptual directness — the ability to systematically construct the appropriate tensor network structure starting from the NQS — is an important advantage of the NQS framework for states with modulated symmetries.
To compare the TPS with the conventional MPS, we employ DMRG to numerically obtain the MPS ground state of the dCSII Hamiltonian. The resulting bond dimension of the MPS tensor $A^{s}$ scales as $\chi=N^{2}$. Meanwhile, the TPS tensors in Eq. (III.2) each carry three virtual indices of dimension NN, so that the local tensor has $N^{4}$ elements — a factor of NN smaller than the $N^{5}$ elements of the MPS tensor with bond dimension $N^{2}$. Thus, at the level of local tensors, the TPS provides a more compact representation of the dCSII ground state.
We note, however, that the compactness of local tensors does not directly translate into computational efficiency for the full network contraction. In the MPS representation, the cost of computing the norm or local observables scales as $O(N^{7}L)$. In the TPS representation, the third virtual index $\gamma_{j}$ connects non-nearest-neighbor sites and introduces loops into the tensor network. Even with an optimized contraction ordering, these loops incur an additional factor of NN, leading to a contraction cost of $O(N^{8}L)$. Thus, despite the smaller local tensor size, exact contraction of the TPS is a factor of NN more expensive than MPS. The advantage of the NQS/TPS construction for the dCSII state is therefore primarily representational — it reveals the natural tensor structure dictated by the modulated symmetries — rather than computational.
IV Kramers-Wannier operation as dipolar Fourier transform and its MPO
Recent investigation has identified another symmetry of the cluster model which is, unlike the other global and modulated symmetries, non-invertible [1, 32, 31, 5, 19, 22, 23, 28, 37]. The ${\mathbb{Z}}_{N}$ Kramers-Wannier (KW) operator implementing such non-invertible symmetry (NIS) is
$K=\sum_{\mathbf{g},\mathbf{g}^{\prime}}\omega^{\sum_{j}(g_{j}-g_{j+1})g^{\prime}_{j}}|\mathbf{g}^{\prime}\rangle\langle\mathbf{g}|,$ (61)
with $\mathbf{g}=\{g_{1},g_{2},\cdots\}$ and $\mathbf{g}^{\prime}=\{g^{\prime}_{1},g^{\prime}_{2},\cdots\}$ being the collection of ${\mathbb{Z}}_{N}$ variables. The non-invertible structures of the cluster state are part of the broader framework of non-invertible symmetries recently studied in quantum field theory and condensed matter systems [3, 32].
Readers who are familiar with the work of [1] will realize that the expression in (61) is the ${\mathbb{Z}}_{N}$ generalization of the $N=2$ KW operator introduced there. In [30], the authors provided the MPO version of the KW operator, which we can also generalize to ${\mathbb{Z}}_{N}$,
$K$ $={\rm Tr}[{\cal K}^{1}\otimes\cdots\otimes{\cal K}^{L}]$
$=\sum_{\mathbf{g}}({\cal K}^{1})_{g_{1},g_{2}}\otimes\cdots\otimes({\cal K}^{L})_{g_{L},g_{1}},$ (62)
where ${\bf g}=\{g_{1},\cdots,g_{L}\}$, and the local MPO tensor ${\cal K}^{j}$ contains the elements which are themselves operators:
$({\cal K}^{j})_{g_{j},g_{j+1}}$ $=\sum_{g^{\prime}_{j}}\omega^{(g_{j}-g_{j+1})g^{\prime}_{j}}|g^{\prime}_{j}\rangle\langle g_{j}|$
$=|\overline{g_{j+1}-g_{j}}\rangle\langle g_{j}|.$ (63)
The overlined state $|\overline{g}\rangle=\sum_{s}\omega^{-sg}|s\rangle$ is the eigenstate of local $X$ operator $X|\overline{g}\rangle=\omega^{g}|\overline{g}\rangle$. One can check that the expression in (63) reduces to the one in [30] by taking $N=2$.
Viewed as the action on a given wavefunction $\Psi({\bf s})$, the KW operator implements the transformation
$\Psi(\mathbf{s})\xrightarrow{K}\Psi^{\prime}({\bf s})=\sum_{\mathbf{g}}\omega^{\sum_{j}g_{j}(s_{j}-s_{j+1})}\Psi(\mathbf{g}).$ (64)
This is reminiscent of the Fourier transformation of the function $\Psi({\bf g})$, except for the subtle difference that the target variable is the difference $s_{j}-s_{j+1}$ rather than $s_{j}$ itself. In fact, this observation leads to very natural interpretation of the KW transformation as the generalization of the ordinary Fourier transformation in which the target variable is not the charge of the target space $s_{j}$, but the dipole in the same space given by $d_{j}\equiv s_{j}-s_{j+1}$. In a sense, the KW transformation is like the dipole Fourier transform while the original Fourier transform would now be the charge Fourier transform. The dipolar Fourier transform is non-invertible because the target variable $d_{j}\equiv s_{j}-s_{j+1}$ has a constraint, $\sum_{j}d_{j}=0$, and covers only $1/N$-th of the target space. The transformed function $\Psi^{\prime}({\bf s})$ must automatically satisfy $\Psi^{\prime}(\{s_{j}\rightarrow s_{j}+1\})=\Psi^{\prime}({\bf s})$, i.e. become an eigenstate of $C=\prod_{j}X_{j}$ with eigenvalue +1.
Further generalization of this observation is to the quadrupole Fourier transformation in which the target variable is $q_{j}=s_{j+1}-2s_{j}+s_{j-1}$:
$\Psi(\mathbf{s})\rightarrow\Psi^{\prime}({\bf s})=\sum_{\mathbf{g}}\omega^{\sum_{j}g_{j}q_{j}}\Psi(\mathbf{g}).$ (65)
Now the target space has two constraints, in the form of the total “charge” and the total “dipole” conservation: $\sum_{j}q_{j}=\sum_{j}jq_{j}=0$  [^1] More precisely, the total dipole moment must be zero modulus NN, assuming that the length of the chain $L$ is divisible by NN.. The transformed function is invariant under both $s_{j}\rightarrow s_{j}+1$ and $s_{j}\rightarrow s_{j}+j$, i.e. resides in the space where both $C$ and $D=\prod_{j}(X_{j})^{j}$ are effectively one. The quadrupole Fourier transformation appears as the KW operator implementing the NIS of the dCSII - see App. F.
The NIS operator specific to the ${\mathbb{Z}}_{N}$ cluster state is
$K_{c}=TK_{1}K_{2},$ (66)
where $T=\sum_{\mathbf{g}}\bigotimes_{j}|g_{j}\rangle_{j+1}\langle g_{j}|_{j}$ gives the translation by one site, and $K_{1}$ and $K_{2}$ are the KW operators acting on the odd and even sublattices,
$K_{1}$ $=\sum_{\mathbf{g}_{1},\mathbf{g}^{\prime}_{1}}\omega^{\sum_{n}(g_{2n-1}-g_{2n+1})g^{\prime}_{2n-1}}|\mathbf{g}^{\prime}_{1}\rangle\langle\mathbf{g}_{1}|,$
$K_{2}$ $=\sum_{\mathbf{g}_{2},\mathbf{g}^{\prime}_{2}}\omega^{\sum_{n}(g_{2n}-g_{2n+2})g^{\prime}_{2n}}|\mathbf{g}^{\prime}_{2}\rangle\langle\mathbf{g}_{2}|,$ (67)
with ${\bf g}_{1}=\{g_{1},g_{3},\cdots g_{L-1}\}$ and ${\bf g}_{2}=\{g_{2},g_{4},\cdots g_{L}\}$, respectively, for even $L$. Explicitly,
$K_{c}=\sum_{\mathbf{g},\mathbf{g}^{\prime}}\omega^{\sum_{j}(g_{j-1}-g_{j+1})g^{\prime}_{j}}|\mathbf{g}^{\prime}\rangle\langle\mathbf{g}|,$ (68)
where ${\bf g},{\bf g}^{\prime}$ span all the qudits of the lattice. Invariance of the CS wavefunction under $K_{c}$ is easily checked:
$\Psi(\mathbf{s})\xrightarrow{K_{c}}\sum_{\mathbf{g}}\omega^{\sum_{j}(s_{j+1}-s_{j-1})g_{j}}\Psi(\mathbf{g})\propto\Psi(\mathbf{s}).$ (69)
One can see that this is dipole Fourier transformation to $d_{j}\equiv s_{j+1}-s_{j-1}$. The new variable satisfies $\sum_{j\in{\rm odd}}d_{j}=\sum_{j\in{\rm even}}d_{j}=0$ over the odd and even sublattices, separately. Functions that emerge from the $K_{c}$ mapping automatically satisfy $C_{1}=C_{2}=1$. For completeness, we list the fusion rules satisfied by the three symmetry generators $\{K_{c},C_{1},C_{2}\}$ of the cluster state [19]:
$C_{1}K_{c}=K_{c}C_{1}=K_{c},\quad C_{2}K_{c}=K_{c}C_{2}=K_{c},$
$K_{c}^{\dagger}K_{c}=K_{c}^{2}=\left(\sum_{m=1}^{N}C_{1}^{m}\right)\left(\sum_{n=1}^{N}C_{2}^{n}\right).$ (70)
Figure 7: $P^{2}$-representation of the $K_{c}$ operator as an MPO.
The MPO of the KW operator $K_{c}$ finds a compact expression in the $P$-representation. Following the derivation in App. E, the matrix elements of $K_{c}$ are
$\langle{\bf s}^{\prime}|K_{c}|{\bf s}\rangle=$ ${\rm Tr}[{\bm{\Omega}}^{*}P^{s_{1}}{\bm{\Omega}}P^{s^{\prime}_{2}}{\bm{\Omega}}^{*}P^{s_{3}}\cdots{\bm{\Omega}}P^{s^{\prime}_{L}}]$
$\times{\rm Tr}[{\bm{\Omega}}P^{s^{\prime}_{1}}{\bm{\Omega}}^{*}P^{s_{2}}{\bm{\Omega}}P^{s^{\prime}_{3}}\cdots{\bm{\Omega}}^{*}P^{s_{L}}].$ (71)
Each trace is in fact a CS wavefunction or its complex conjugate, written in a sort of doubled Hilbert space of $\bf s$ and $\bf s^{\prime}$. Based on this observation and the fact that each CS wavefunction is invariant under the $C_{1}$ and $C_{2}$ symmetries, it follows that $K_{c}$ is also invariant under each of them. This explains the first two fusion rules in (70). Invoking the theorem in linear algebra, the product of two traces can be combined as a single trace:
$\langle{\bf s}^{\prime}|K_{c}|{\bf s}\rangle=$
${\rm Tr}[({\bm{\Omega}}^{*}P^{s_{1}}\otimes{\bm{\Omega}}P^{s^{\prime}_{1}})({\bm{\Omega}}P^{s^{\prime}_{2}}\otimes{\bm{\Omega}}^{*}P^{s_{2}})\cdots({\bm{\Omega}}P^{s^{\prime}_{L}}\otimes{\bm{\Omega}}^{*}P^{s_{L}})].$ (72)
Since the MPO now involves a pair of projectors at each site, we refer to it as the $P^{2}$-representation of the MPO for the KW operator. Figure 7 illustrates the MPO of the KW operator $K_{c}$ graphically.
We can write the KW operator itself, rather than its matrix elements, as an MPO:
$K_{c}$ $={\rm Tr}[{\cal K}^{1}_{c}\otimes\cdots\otimes{\cal K}^{L}_{c}]$
$=\sum_{\mathbf{\bm{\alpha}}}({\cal K}^{1}_{c})_{\alpha_{L},\alpha_{1}}\otimes\cdots\otimes({\cal K}^{L}_{c})_{\alpha_{L-1},\alpha_{L}},$ (73)
where ${\bm{\alpha}}=\{\alpha_{1},\cdots,\alpha_{L}\}$ and each bond index $\alpha_{j}\equiv(g_{j},h_{j})$ consists of a pair of ${\mathbb{Z}}_{N}$ numbers. Each element of the matrix ${\cal K}^{j}_{c}$ is an operator acting on the qudit at the site $j$. To match the decomposition (72) of $K_{c}$, we invoke the identification
$\langle s^{\prime}_{j}|({\cal K}^{j}_{c})_{\alpha_{j-1},\alpha_{j}}|s_{j}\rangle=$ $({\bm{\Omega}}^{*}P^{s_{j}}\otimes{\bm{\Omega}}^{*}P^{s^{\prime}_{j}})_{\alpha_{j-1},\alpha_{j}}$
$=$ $({\bm{\Omega}}^{*}P^{s_{j}})_{g_{j-1},g_{j}}({\bm{\Omega}}P^{s^{\prime}_{j}})_{h_{j-1},h_{j}}$
$=$ $\omega^{-g_{j-1}g_{j}}\delta_{g_{j},s_{j}}\omega^{h_{j-1}h_{j}}\delta_{h_{j},s^{\prime}_{j}}$ (74)
for odd $j$ and
$\langle s^{\prime}_{j}|({\cal K}^{j}_{c})_{\alpha_{j-1},\alpha_{j}}|s_{j}\rangle=$ $({\bm{\Omega}}P^{s^{\prime}_{j}}\otimes{\bm{\Omega}}^{*}P^{s_{j}})_{\alpha_{j+1},\alpha_{j}}$
$=$ $({\bm{\Omega}}P^{s^{\prime}_{j}})_{g_{j-1},g_{j}}({\bm{\Omega}}^{*}P^{s_{j}})_{h_{j-1},h_{j}}$
$=$ $\omega^{g_{j-1}g_{j}}\delta_{g_{j},s^{\prime}_{j}}\omega^{-h_{j-1}h_{j}}\delta_{h_{j},s_{j}}$ (75)
for even $j$. This leads to the definition
$({\cal K}^{j}_{c})_{\alpha_{j-1},\alpha_{j}}=\begin{cases}\omega^{h_{j-1}h_{j}-g_{j-1}g_{j}}|h_{j}\rangle\langle g_{j}|,&j~{\rm odd},\\ \omega^{g_{j-1}g_{j}-h_{j-1}h_{j}}|g_{j}\rangle\langle h_{j}|,&j~{\rm even}.\end{cases}$ (76)
It turns out that the MPO representation of the KW operator comes in many different forms. A collection of different MPOs for $K_{c}$ is given in App. E. The KW NIS operator for the dCSII state is discussed in App. F for completeness.
Although the discussion of KW symmetry lies somewhat outside the main focus of the paper, we include it here for three reasons. First, it extends the previously known ${\mathbb{Z}}_{2}$ MPO representation of the KW operator [30] to ${\mathbb{Z}}_{N}$. Second, it illustrates the utility of the $P$-representation not only for expressing cluster states themselves, but also for constructing NIS operators as MPO. Finally, we arrive at an intuitively appealing interpretation of the KW transformation as the dipolar version of the discrete Fourier transform with the obvious restriction on the target space giving clear intuition to the non-invertibility of the mapping.
V Summary and discussion
Neural quantum states (NQS) have recently emerged as a powerful framework for encoding many-body quantum ground states, motivated in large part by advances in AI community. In this work, we revisited the $Z_{N}$ cluster and dipolar cluster states as paradigmatic one-dimensional SPT states protected by spatially uniform and modulated symmetries, and analyzed them with the view of NQS in mind. For all the cluster models analyzed in this work, we arrived at a compact description called the $P$-representation, in which the physical degrees of freedom are encoded solely through a local projector $P^{s}$. The interaction matrix $\bm{\Omega}$ then encodes the patterns by which the local spins are connected to other spins. When each projector carries two virtual legs, the $P$-representation becomes equivalent to the familiar MPS description of the cluster state. When three virtual legs are present, as in the dCSII state, the resulting construction realizes a tensor product state instead. Within this framework, standard SPT features—most notably the action of symmetries on virtual degrees of freedom—admit particularly simple graphical proofs as demonstrated for all three cluster models.
The NQS representation is obtained by factorizing the interaction matrix in the $P$-representation into a pair of weight matrices via a singular-value–like decomposition. For the cluster states studied here, this decomposition can be carried out analytically, yielding closed-form expressions for the NQS weights in all three cases. A useful by-product of our analysis is an explicit matrix-product-operator (MPO) form of the Kramers–Wannier duality operator, which, together with the NQS construction of the cluster states, generalizes earlier ${\mathbb{Z}}_{2}$ results to arbitrary ${\mathbb{Z}}_{N}$.
Perturbations of the cluster Hamiltonians will in general deform the corresponding ground states. In the present language, such deformations can be viewed as renormalizations of the interaction matrix $\bm{\Omega}$, or the weight matrix $\bf W$}, in analogy to how a Slater determinant is modified by a Jastrow factor or how fixed-point tensors are renormalized away from exact SPT fixed points. It is plausible that these renormalizations can be implemented efficiently using modern AI-based techniques for learning many-body wavefunctions [38], or that the nonlinearity structure of the weight functions is tied to the entanglement content of the state [25]. Exploring these directions—both for perturbed cluster models and for more general modulated SPT states—remains an interesting avenue for future research. Representing a two-dimensional topological state such as the toric code and SPT states with ${\mathbb{Z}}_{N}$ physical degrees of freedom is another exciting avenue of future research.
Acknowledgements.
JHH was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (Grant No. 2023R1A2C1002644 and No. RS-2024-00410027). He thanks Ömer M. Aksoy, Zhian Jia, Nisarga Paul, Shu-Heng Shao, Xiao-Gang Wen, Yizhi You for informative discussion. H.-Y.L was supported by the Basic Science Research Program through the National Research Foundation of Korea funded by the Ministry of Science and ICT [Grant No. RS-2023-00220471, RS-2025-16064392].
Appendix A Proof of symmetry fractionalization in the cluster state
Suppose we performed the $C_{1}^{\dagger}$ transformation on the open-chain cluster ground state (11), to get
$C_{1}^{\dagger}|\Psi\rangle=$ $\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|A^{s_{2}}\cdots A^{s_{L}}|s_{L}\rangle|s_{1}-1,s_{2},s_{3}-1,\dots\rangle$
$=$ $X_{1}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|A^{s_{2}}A^{s_{3}+1}A^{s_{4}}\cdots A^{s_{L-1}+1}A^{s_{L}}|s_{L}\rangle|\mathbf{s}\rangle.$ (77)
Applying the identities $A^{s_{2j-1}+1}=Z^{\dagger}A^{s_{2j-1}}X^{\dagger}$ and $X^{\dagger}A^{s_{2j}}Z^{\dagger}=A^{s_{2j}}$ successively, we are left with
$C_{1}^{\dagger}|\Psi\rangle=$ $X_{1}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|A^{s_{2}}Z^{\dagger}A^{s_{3}}A^{s_{4}}\cdots A^{s_{L}}Z|s_{L}\rangle|{\bf s}\rangle$
$=$ $X_{1}^{\dagger}\omega^{-s_{2}+s_{L}}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|A^{s_{2}}A^{s_{3}}A^{s_{4}}\cdots A^{s_{L}}|s_{L}\rangle|{\bf s}\rangle$
$=$ $X_{1}^{\dagger}Z_{2}^{\dagger}Z_{L}|\Psi\rangle.$ (78)
Similarly, performing $C_{2}^{\dagger}$ on an open-chain state results in
$C_{2}^{\dagger}|\Psi\rangle=$ $\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|A^{s_{2}}\cdots A^{s_{L}}|s_{L}\rangle|s_{1},s_{2}-1,\dots,s_{L}-1\rangle$
$=$ $X_{L}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|A^{s_{2}+1}A^{s_{3}}A^{s_{4}+1}\cdots A^{s_{L}}|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $X_{L}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|ZA^{s_{2}}A^{s_{3}}A^{s_{4}}\cdots A^{s_{L-1}}Z^{\dagger}A^{s_{L}}|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $X_{L}^{\dagger}\omega^{s_{1}-s_{L-1}}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|A^{s_{2}}A^{s_{3}}A^{s_{4}}\cdots A^{s_{L}}|s_{L}\rangle|{\bf s}\rangle$
$=$ $Z_{1}Z_{L-1}^{\dagger}X_{L}^{\dagger}|\Psi\rangle.$ (79)
The proof of symmetry fractionalization proceeds in a similar fashion for the P-representation. Acting on the open-chain ground state (29) with $C_{1}$ yields
$C_{1}^{\dagger}|\Psi\rangle=$ $X_{1}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|\bm{\Omega}P^{s_{2}}\bm{\Omega}^{*}(XP^{s_{3}}X^{\dagger})\cdots$
$\mkern 60.0mu\cdots(XP^{s_{L-1}}X^{\dagger})\bm{\Omega}|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $X_{1}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|\bm{\Omega}Z^{\dagger}(ZP^{s_{2}}Z^{\dagger})\cdots$
$\mkern 60.0mu\cdots\bm{\Omega}(ZP^{s_{L-2}}Z^{\dagger})\bm{\Omega}^{*}P^{s_{L-1}}\bm{\Omega}Z|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $X_{1}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|\bm{\Omega}Z^{\dagger}P^{s_{2}}\bm{\Omega}^{*}P^{s_{3}}\bm{\Omega}\cdots P^{s_{L-1}}\bm{\Omega}Z|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $X_{1}^{\dagger}\omega^{-s_{2}+s_{L}}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|\bm{\Omega}P^{s_{2}}\bm{\Omega}^{*}P^{s_{3}}\bm{\Omega}\cdots P^{s_{L-1}}\bm{\Omega}|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $X_{1}^{\dagger}Z_{2}^{\dagger}Z_{L}|\Psi\rangle.$ (80)
A similar exercise gives
$C_{2}^{\dagger}|\Psi\rangle=$ $X_{L}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|\bm{\Omega}(XP^{s_{2}}X^{\dagger})\cdots$
$\mkern 60.0mu\cdots(XP^{L-2}X^{\dagger})\bm{\Omega}^{*}P^{s_{L-1}}\bm{\Omega}|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $X_{L}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|Z\bm{\Omega}P^{s_{2}}\bm{\Omega}^{*}(Z^{\dagger}P^{s_{3}}Z)\bm{\Omega}\cdots$
$\mkern 60.0mu\cdots(Z^{\dagger}P^{s_{L-1}}Z)Z^{\dagger}\bm{\Omega}|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $X_{L}^{\dagger}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|Z\bm{\Omega}P^{s_{2}}\bm{\Omega}^{*}P^{s_{3}}\bm{\Omega}\cdots P^{s_{L-1}}Z^{\dagger}\bm{\Omega}|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $X_{L}^{\dagger}\omega^{s_{1}-s_{L-1}}\mkern-12.0mu\sum_{s_{2},\dots,s_{L-1}}\mkern-12.0mu\langle s_{1}|\bm{\Omega}P^{s_{2}}\bm{\Omega}^{*}P^{s_{3}}\bm{\Omega}\cdots P^{s_{L-1}}\bm{\Omega}|s_{L}\rangle|\mathbf{s}\rangle$
$=$ $Z_{1}Z_{L-1}^{\dagger}X_{L}^{\dagger}|\Psi\rangle.$ (81)
Appendix B Derivation of weight matrix $W(s,h)$ for the ${\mathbb{Z}}_{N}$ cluster state
In this section we introduce a general scheme to construct the function $W$ satisfying $\sum_{h}W(x,h)W(y,h)=\omega^{pxy}$. Starting from the ansatz
$W(x,h)=\frac{1}{\sqrt{\kappa}}\omega^{ah^{2}+bx^{2}+cxh}\,$, (82)
we directly get
$\sum_{h}W(x,h)W(y,h)=\frac{1}{\kappa}\omega^{b(x^{2}+y^{2})}\sum_{h}\omega^{2ah^{2}+c(x+y)h}\,.$ (83)
For some integer $r$ satisfying the equation
$c(x+y)-4ar=0\mod N\,$, (84)
we can complete the square and write
$\sum_{h}W(x,h)W(y,h)=\frac{1}{\kappa}\omega^{-2ar^{2}+b(x^{2}+y^{2})}\sum_{h}\omega^{2a(h+r)^{2}}\,.$ (85)
The conditions for the periodicity of the summand in Eq. (85) are
$2a$ $\in\mathbb{Z}\quad\text{for odd }N\,$,
$4a$ $\in\mathbb{Z}\quad\text{for even }N\,.$ (86)
Employing the constant $a$ that ensures periodicity of the summand, Eq. (85) becomes
$\sum_{h}W(x,h)W(y,h)=\frac{1}{\kappa}\omega^{-2ar^{2}+b(x^{2}+y^{2})}\sum_{h}\omega^{2ah^{2}}\,.$ (87)
Then it is natural to define the normalization factor $\kappa$ that supposed to be nonzero as
$\kappa=\sum_{h}\omega^{2ah^{2}}\neq 0\,.$ (88)
When we further invoke the equation
$2ar^{2}=b(x+y)^{2}\mod N\,$, (89)
we finally obtain
$\sum_{h}W(x,h)W(y,h)=\omega^{-2bxy}\,.$ (90)
To make the triplet $(a,b,c)$ is constant with respect to the arguments $x$ and $y$, we set
$r=(x+y)k\,$, (91)
for some integer $k$. Plugging this into Eq. (84) and Eq. (89), we have
$c$ $=4ak\mod N\,$,
$b$ $=2ak^{2}\mod N\,.$ (92)
Let us assume $k=\pm 1$ for a simple solution and the conditions Eq. (B) and Eq. (88) are satisfied. Then the triplet
$(a,b,c)=\left(-\frac{p}{4},-\frac{p}{2},\pm p\right)$ (93)
gives
$\sum_{h}W(x,h)W(y,h)=\omega^{pxy}\,$, (94)
for $p\in\mathbb{Z}_{N}$, and $W$ constructs an RBM representation of the SPT states associated with $H^{2}(\mathbb{Z}_{N}\times\mathbb{Z}_{N},U(1))$.
Appendix C Proof of symmetry invariance for MPS of the cluster state
Next we use the MPS representation in (23) to prove the invariance, which turns out to be much more involved than the above proof made in the NQS representation. To this end, we need to know how $\bf W$} transforms with respect to $X$ and $Z$, which are captured by several identities
$\mathbf{W}^{\dagger}X$ $=\omega^{\frac{1}{2}}Z\mathbf{W}^{\dagger}Z\,$,
$X^{\dagger}\mathbf{W}$ $=\omega^{-\frac{1}{2}}Z^{-1}\mathbf{W}Z^{-1}\,$,
$\mathbf{W}^{t}X$ $=\omega^{-\frac{1}{2}}Z^{-1}\mathbf{W}^{t}Z^{-1}\,$,
$X^{\dagger}\mathbf{W}^{*}$ $=\omega^{\frac{1}{2}}Z\mathbf{W}^{*}Z.$ (95)
With these identities the MPSs transform as
$A^{s_{2j+1}+1}$ $=ZA^{s_{2j+1}}Z^{-1},$
$A^{s_{2j}+1}$ $=Z^{-1}A^{s_{2j}}Z.$ (96)
From the expression of the MPS (23) and the identities (C) we show that applying $C_{1}^{\dagger}$ to the CS wavefunction changes it to
$\Psi(\mathbf{s})$ $\xrightarrow{C_{1}^{\dagger}}\mathrm{Tr}\left[\mathbf{W}^{\dagger}(XP^{s_{1}}X^{\dagger})\mathbf{W}\cdot\mathbf{W}^{t}P^{s_{2}}\mathbf{W}^{*}\cdots\mathbf{W}^{t}P^{s_{L}}\mathbf{W}^{*}\right]$
$=\mathrm{Tr}\!\left[(Z\mathbf{W}^{\dagger}P^{s_{1}}\mathbf{W}Z^{-1})\cdot\mathbf{W}^{t}P^{s_{2}}\mathbf{W}^{*}\!\cdots\!\mathbf{W}^{t}P^{s_{L}}\mathbf{W}^{*}\right].$ (97)
It remains to show how, despite these transformations on the MPS matrix, the overall wavefunction stays invariant. Let us first re-organize the expression in Eq. (C) as
$\mathrm{Tr}\left[(\mathbf{W}^{*}Z\mathbf{W}^{\dagger})P^{s_{1}}(\mathbf{W}Z^{-1}\mathbf{W}^{t})P^{s_{2}}(\mathbf{W}^{*}Z\mathbf{W}^{\dagger})\cdots{\bf W}^{t}P^{s_{L}}\right]$ (98)
where the cyclicity of the trace has been used to move $\bf W$$^{*}$ at the end of the trace to the beginning of it. To make further progress we use following identities:
$\mathbf{W}Z^{-1}\mathbf{W}^{t}$ $=\omega^{-\frac{1}{2}}X^{\dagger}\mathbf{W}\mathbf{W}^{t}X=\omega^{-\frac{1}{2}}X^{\dagger}\bm{\Omega}X$
$\mathbf{W}^{*}Z\mathbf{W}^{\dagger}$ $=\omega^{\frac{1}{2}}X^{\dagger}\mathbf{W}^{*}\mathbf{W}^{\dagger}X=\omega^{\frac{1}{2}}X^{\dagger}\bm{\Omega}^{*}X$ (99)
Plugging (C) into (98) gives
$\mathrm{Tr}\left[(X^{\dagger}\bm{\Omega}^{*}X)P^{s_{1}}(X^{\dagger}\bm{\Omega}X)P^{s_{2}}(X^{\dagger}\bm{\Omega}^{*}X)\cdots(X^{\dagger}\bm{\Omega}X)P^{s_{L}}\right]$
$\!\!=\mathrm{Tr}\left[(\bm{\Omega}^{*}Z^{\dagger}X)P^{s_{1}}(\bm{\Omega}ZX)P^{s_{2}}(\bm{\Omega}^{*}Z^{\dagger}X)\cdots(\bm{\Omega}ZX)P^{s_{L}}\right]$
$\!\!=\mathrm{Tr}\left[(\bm{\Omega}^{*}\omega^{-1}XZ^{\dagger})P^{s_{1}}(\bm{\Omega}\omega XZ)P^{s_{2}}(\bm{\Omega}^{*}Z^{\dagger}X)\cdots(\bm{\Omega}ZX)P^{s_{L}}\right]$
$\!\!=\mathrm{Tr}\left[(\bm{\Omega}^{*}XZ^{\dagger})P^{s_{1}}(\bm{\Omega}XZ)P^{s_{2}}(\bm{\Omega}^{*}XZ^{\dagger})\cdots(\bm{\Omega}XZ)P^{s_{L}}\right]$
$\!\!=\mathrm{Tr}\left[(Z^{\dagger}\bm{\Omega}^{*}Z^{\dagger})P^{s_{1}}(Z\bm{\Omega}Z)P^{s_{2}}(Z^{\dagger}\bm{\Omega}^{*}Z^{\dagger})\cdots(Z\bm{\Omega}Z)P^{s_{L}}\right]$
$\!\!=\mathrm{Tr}\left[\bm{\Omega}^{*}(Z^{\dagger}P^{s_{1}}Z)\bm{\Omega}(ZP^{s_{2}}Z^{\dagger})\bm{\Omega}^{*}Z^{\dagger}\cdots Z\bm{\Omega}(ZP^{s_{L}}Z^{\dagger})\right]$
$\!\!=\mathrm{Tr}\left[\bm{\Omega}^{*}P^{s_{1}}\bm{\Omega}P^{s_{2}}\bm{\Omega}^{*}\cdots\bm{\Omega}P^{s_{L}}\right]=\Psi(\mathbf{s})$ (100)
verifying the invariance of the MPS wavefunction under $C_{1}$. A similar exercise establishes the invariance under $C_{2}$.
The identities of the matrix $\mathbf{W}$ and the $\mathbb{Z}_{N}$ Pauli matrices can be obtained by algebraic calculation of thier matrix elements. The first line of Eq. (C) can be explicitly shown by
$[\mathbf{W}^{\dagger}X]_{\alpha\beta}$ $\equiv\langle\alpha|\mathbf{W}^{\dagger}X|\beta\rangle=\langle\alpha|\mathbf{W}^{\dagger}|\beta+1\rangle$
$=\frac{1}{\sqrt{\kappa}}\omega^{-(a\alpha^{2}+b(\beta+1)^{2}+c\alpha(\beta+1))}$
$=\frac{1}{\sqrt{\kappa}}\omega^{-(a\alpha^{2}+b\beta^{2}+c\alpha\beta)}\omega^{-2b\beta}\omega^{-c\alpha}\omega^{-b}$
$=\langle\alpha|\omega^{-b}Z^{-c}\mathbf{W}^{\dagger}Z^{-2b}|\beta\rangle\,.$ (101)
The second line is computed by
$[X^{\dagger}\mathbf{W}]_{\alpha\beta}$ $\equiv\langle\alpha|X^{\dagger}\mathbf{W}|\beta\rangle=\langle\alpha+1|\mathbf{W}|\beta\rangle$
$=\frac{1}{\sqrt{\kappa}}\omega^{a\beta^{2}+b(\alpha+1)^{2}+c(\alpha+1)\beta}$
$=\frac{1}{\sqrt{\kappa}}\omega^{a\beta^{2}+b\alpha^{2}+c\alpha\beta+2b\alpha+b+c\beta}$
$=\langle\alpha|\omega^{b}Z^{2b}\mathbf{W}Z^{c}|\beta\rangle$ (102)
We now explicitly calculate $\mathbf{W}Z^{c}\mathbf{W}^{t}$ (C). Using the coordinate representations $Z]_{\alpha\beta}=\delta_{\alpha\beta}\omega^{\alpha}$ and $Z^{c}]_{\alpha\beta}=\delta_{\alpha\beta}\omega^{c\alpha}$, we obtain
$[\mathbf{W}Z^{c}\mathbf{W}^{t}]_{\alpha\eta}$ $=\sum_{\beta\gamma}[\mathbf{W}]_{\alpha\beta}[Z^{c}]_{\beta\gamma}[\mathbf{W}^{t}]_{\gamma\eta}=\frac{1}{\kappa}\omega^{b(\alpha^{2}+\eta^{2})}\sum_{\beta}\omega^{2a\beta^{2}+c\beta(\alpha+\eta+1)}$
$=\omega^{b(\alpha^{2}+\eta^{2})-b\left(\alpha+\eta+1\right)^{2}}=\omega^{-2b(\alpha+1)(\eta+1)}\omega^{b}$
$=\omega^{b}\langle\alpha|X^{\dagger}\mathbf{W}\mathbf{W}^{t}X|\eta\rangle\,.$ (103)
Here we use the relation (B). Similarly, $\mathbf{W}^{*}Z^{-c}\mathbf{W}^{\dagger}$ gives
$[\mathbf{W}^{*}Z^{-c}\mathbf{W}^{\dagger}]_{\alpha\eta}$ $=\sum_{\beta\gamma}[\mathbf{W}^{*}]_{\alpha\beta}[Z^{-c}]_{\beta\gamma}[\mathbf{W}^{\dagger}]_{\gamma\eta}=\frac{1}{\kappa^{*}}\omega^{-b(\alpha^{2}+\eta^{2})}\sum_{\beta}\omega^{-2a\beta^{2}-c\beta(\alpha+\eta+1)}$
$=\omega^{-b(\alpha^{2}+\eta^{2})+b\left(\alpha+\eta+1\right)^{2}}=\omega^{2b(\alpha+1)(\eta+1)}\omega^{-b}$
$=\omega^{-b}\langle\alpha|X^{\dagger}\mathbf{W}^{*}\mathbf{W}^{\dagger}X|\eta\rangle\,.$ (104)
Appendix D Proof of symmetry for dCSII
We prove the NQS representation of the dCSII wavefunction is invariant under $\{C_{1},C_{2},D_{1},D_{2}\}$. Invoking the relation (III.2), the wavefunction under $C_{1}^{\dagger}$ is written as
$\Psi(\mathbf{s})\xrightarrow[]{C^{\dagger}_{1}}\prod_{j}\sum_{\alpha_{j}\beta_{j}\gamma_{j}}$ $W(s_{2j-1}+1,\alpha_{j})W(s_{2j},\alpha_{j})\tilde{W}(s_{2j},\beta_{j})\tilde{W}(s_{2j+1}+1,\beta_{j})W(s_{2j},\gamma_{j})W(s_{2j+3}+1,\gamma_{j})$
$=\prod_{j}\sum_{\alpha_{j}\beta_{j}\gamma_{j}}$ $\bigg[W(s_{2j-1},\alpha_{j})W(s_{2j},\alpha_{j})\tilde{W}(s_{2j},\beta_{j})\tilde{W}(s_{2j+1},\beta_{j})W(s_{2j},\gamma_{j})W(s_{2j+3},\gamma_{j})$
$\times\omega^{2b(s_{2j-1}+s_{2j+3}-2s_{2j+1})}\omega^{c(\alpha_{j}+\gamma_{j})}\omega^{\tilde{c}\beta_{j}}\bigg]$
$=\prod_{j}\sum_{\alpha_{j}\beta_{j}\gamma_{j}}$ $W(s_{2j-1},\alpha_{j})W(s_{2j},\alpha_{j})\tilde{W}(s_{2j},\beta_{j})\tilde{W}(s_{2j+1},\beta_{j})W(s_{2j},\gamma_{j})W(s_{2j+3},\gamma_{j})\omega^{c(\alpha_{j}+\gamma_{j})}\omega^{\tilde{c}\beta_{j}}\,.$ (105)
The summation over the hidden variable $\alpha_{j}$ is explicitly given by
$\sum_{\alpha_{j}}W(s_{2j-1},\alpha_{j})W(s_{2j},\alpha_{j})\omega^{c\alpha_{j}}$ $=\omega^{-b}\omega^{-2b(s_{2j-1}s_{2j})}\omega^{-2b(s_{2j+1}+s_{2j})}\,.$ (106)
Similar to this we can explicitly sum over all hidden variables, and the wavefunction (D) is written as
$\prod_{j}\omega^{-2b(s_{2j-1}s_{2j})}\omega^{-2\tilde{b}(s_{2j}s_{2j+1})}\omega^{-2b(s_{2j}s_{2j+3})}\omega^{-2b(s_{2j+1}+2s_{2j}+s_{2j+3})}\omega^{-2\tilde{b}(s_{2j}+s_{2j+1})}\,.$ (107)
Since the constants $b$ and $\tilde{b}$ satisfy
$\omega^{-2bxy}=\omega^{xy}\,,\quad\omega^{-2\tilde{b}xy}=\omega^{-2xy}\,$, (108)
the Eq. (107) recovers the original wavefunction
$\prod_{j}\omega^{-2b(s_{2j-1}s_{2j})}\omega^{-2\tilde{b}(s_{2j}s_{2j+1})}\omega^{-2b(s_{2j}s_{2j+3})}\omega^{-2b(s_{2j+1}+2s_{2j}+s_{2j+3})}\omega^{-2\tilde{b}(s_{2j}+s_{2j+1})}$
$=\prod_{j}\omega^{s_{2j-1}s_{2j}}\omega^{-2s_{2j}s_{2j+1}}\omega^{s_{2j}s_{2j+3}}=\Psi(\mathbf{s})\,.\,$ (109)
Proof for $C_{2}$ is similar.
Now let us consider the dipole symmetry of the dipolar cluster state. Utilizing the relation (III.2) we show that the wavefunction is invariant under the operation $D_{1}^{\dagger}$
$\Psi(\mathbf{s})\xrightarrow[]{D^{\dagger}_{1}}$ $\prod_{j}\sum_{\alpha_{j}\beta_{j}\gamma_{j}}W(s_{2j-1}+j,\alpha_{j})W(s_{2j},\alpha_{j})\tilde{W}(s_{2j},\beta_{j})\tilde{W}(s_{2j+1}+j+1,\beta_{j})W(s_{2j},\gamma_{j})W(s_{2j+3}+j+2,\gamma_{j})$
$=$ $\prod_{j}\sum_{\alpha_{j}\beta_{j}\gamma_{j}}\bigg[W(s_{2j-1},\alpha_{j})W(s_{2j},\alpha_{j})\tilde{W}(s_{2j},\beta_{j})\tilde{W}(s_{2j+1},\beta_{j})W(s_{2j},\gamma_{j})W(s_{2j+3},\gamma_{j})$
$\times\omega^{2bjs_{2j-1}+cj\alpha_{j}+bj^{2}}\omega^{2\tilde{b}(j+1)s_{2j+1}+\tilde{c}(j+1)\beta_{j}+\tilde{b}(j+1)^{2}}\omega^{2b(j+2)s_{2j+3}+c(j+2)\gamma_{j}+b(j+2)^{2}}\bigg]$
$=$ $\prod_{j}\sum_{\alpha_{j}\beta_{j}\gamma_{j}}\bigg[W(s_{2j-1},\alpha_{j})W(s_{2j},\alpha_{j})\tilde{W}(s_{2j},\beta_{j})\tilde{W}(s_{2j+1},\beta_{j})W(s_{2j},\gamma_{j})W(s_{2j+3},\gamma_{j})\omega^{c(j\alpha_{j}+(j+2)\gamma_{j}-2(j+1)\beta_{j})}\omega^{2b}\bigg]$
$=$ $\prod_{j}\omega^{s_{2j-1}s_{2j}}\omega^{-2s_{2j}s_{2j+1}}\omega^{s_{2j}s_{2j+3}}=\Psi(\mathbf{s})\,.$ (110)
Again a similar proof applies for symmetry under $D_{2}$.
We can also show that the $P$-representation (43) is invariant under $\{C_{1},C_{2},D_{1},D_{2}\}$. Let us see how the projector representation transforms under $C_{2}^{\dagger}$:
$\Psi(\mathbf{s})$ $\xrightarrow[]{C_{2}^{\dagger}}\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}+1}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\tilde{\Omega}_{\beta^{\prime}_{j}\beta_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$
$=$ $\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{(\alpha^{\prime}_{j}+1)\alpha_{j}}\tilde{\Omega}_{(\beta^{\prime}_{j}+1)\beta_{j}}\Omega_{(\gamma^{\prime}_{j}+1)\gamma_{j}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$
$=$ $\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\tilde{\Omega}_{\beta^{\prime}_{j}\beta_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}\omega^{\alpha_{j}-2\beta_{j}+\gamma_{j}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$
$=$ $\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\tilde{\Omega}_{\beta^{\prime}_{j}\beta_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}\omega^{\alpha_{j+1}-2\beta_{j}+\gamma_{j-1}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$
$=$ $\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\tilde{\Omega}_{\beta^{\prime}_{j}\beta_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}=\Psi(\mathbf{s})\,.$ (111)
In going from the first to the second line, we relabel the summation indices as $\alpha^{\prime}_{j}\to\alpha^{\prime}_{j}+1$, $\beta^{\prime}_{j}\to\beta^{\prime}_{j}+1$, and $\gamma^{\prime}_{j}\to\gamma^{\prime}_{j}+1$. In going from the second to the third line, we use the definition of $\Omega$ to extract the factor $\omega^{\alpha_{j}-2\beta_{j}+\gamma_{j}}$. Finally, from the third to the fourth line, we use the property $\prod_{j}\omega^{\alpha_{j}}=\prod_{j}\omega^{\alpha_{j+1}}$.
Similarly, the transformation of the wavefunction under $D_{2}^{\dagger}$ is cast as
$\Psi(\mathbf{s})$ $\xrightarrow[]{D_{2}^{\dagger}}\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}+j}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\tilde{\Omega}_{\beta^{\prime}_{j}\beta_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$
$=$ $\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{(\alpha^{\prime}_{j}+j)\alpha_{j}}\tilde{\Omega}_{(\beta^{\prime}_{j}+j)\beta_{j}}\Omega_{(\gamma^{\prime}_{j}+j)\gamma_{j}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$
$=$ $\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\tilde{\Omega}_{\beta^{\prime}_{j}\beta_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}\omega^{j(\alpha_{j}-2\beta_{j}+\gamma_{j})}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$
$=$ $\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\tilde{\Omega}_{\beta^{\prime}_{j}\beta_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}\omega^{(j+1)\alpha_{j+1}-2j\beta_{j}+(j-1)\gamma_{j-1}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}$
$=$ $\sum_{\bm{\alpha}\bm{\beta}\bm{\gamma}}\sum_{\bm{\alpha}^{\prime}\bm{\beta}^{\prime}\bm{\gamma}^{\prime}}\prod_{j}P^{s_{2j}}_{\alpha^{\prime}_{j}\beta^{\prime}_{j}\gamma^{\prime}_{j}}\Omega_{\alpha^{\prime}_{j}\alpha_{j}}\tilde{\Omega}_{\beta^{\prime}_{j}\beta_{j}}\Omega_{\gamma^{\prime}_{j}\gamma_{j}}P^{s_{2j+1}}_{\alpha_{j+1}\beta_{j}\gamma_{j-1}}=\Psi(\mathbf{s})\,.$ (112)
Equation (D) follows by steps entirely analogous to those used in deriving Eq. (D).
Appendix E Kramers-Wannier transformation as MPO
We first note that the elements of $K_{c}$ are
$\langle{\bf s}^{\prime}|K_{c}|{\bf s}\rangle$ $=\omega^{\sum_{j}(s_{j-1}-s_{j+1})s^{\prime}_{j}}$
$=\omega^{(s_{1}-s_{3})s^{\prime}_{2}+(s_{3}-s_{5})s^{\prime}_{4}+\cdots+(s_{L-1}-s_{1})s^{\prime}_{L}}$
$\times\omega^{(s_{L}-s_{2})s^{\prime}_{1}+(s_{2}-s_{4})s^{\prime}_{3}+\cdots+(s_{L-2}-s_{L})s^{\prime}_{L-1}}.$ (113)
The expression in the first line of (E) has been deliberately split into two parts over the next two lines, each line corresponding to a connected curve going $s_{1}\rightarrow s^{\prime}_{2}\rightarrow s_{3}\rightarrow\cdots$ or $s^{\prime}_{1}\rightarrow s_{2}\rightarrow s^{\prime}_{3}\rightarrow\cdots$ in Fig. 7. Each curve can be expressed as a trace:
$\omega^{(s_{1}-s_{3})s^{\prime}_{2}+(s_{3}-s_{5})s^{\prime}_{4}+\cdots+(s_{L-1}-s_{1})s^{\prime}_{L}}$ $={\rm Tr}[{\bm{\Omega}}^{*}P^{s_{1}}{\bm{\Omega}}P^{s^{\prime}_{2}}{\bm{\Omega}}^{*}P^{s_{3}}\cdots{\bm{\Omega}}P^{s^{\prime}_{L}}]$
$\omega^{(s_{L}-s_{2})s^{\prime}_{1}+(s_{2}-s_{4})s^{\prime}_{3}+\cdots+(s_{L-2}-s_{L})s^{\prime}_{L-1}}$ $={\rm Tr}[{\bm{\Omega}}P^{s^{\prime}_{1}}{\bm{\Omega}}^{*}P^{s_{2}}{\bm{\Omega}}P^{s^{\prime}_{3}}\cdots{\bm{\Omega}}^{*}P^{s_{L}}].$ (114)
Combining the two expressions gives
$\langle{\bf s}^{\prime}|K_{c}|{\bf s}\rangle$ $={\rm Tr}[{\bm{\Omega}}^{*}P^{s_{1}}{\bm{\Omega}}P^{s^{\prime}_{2}}{\bm{\Omega}}^{*}P^{s_{3}}\cdots{\bm{\Omega}}P^{s^{\prime}_{L}}]\times{\rm Tr}[{\bm{\Omega}}P^{s^{\prime}_{1}}{\bm{\Omega}}^{*}P^{s_{2}}{\bm{\Omega}}P^{s^{\prime}_{3}}\cdots{\bm{\Omega}}^{*}P^{s_{L}}].$ (115)
Invoking the theorem
${\rm Tr}[A_{1}A_{2}\cdots A_{L}]{\rm Tr}[B_{1}B_{2}\cdots B_{L}]={\rm Tr}[(A_{1}\otimes B_{1})(A_{2}\otimes B_{2})\cdots(A_{L}\otimes B_{L})]$
allows us to reduce the two-trace expression in Eq. (115) to a single trace:
$\langle{\bf s}^{\prime}|K_{c}|{\bf s}\rangle$ $={\rm Tr}[({\bm{\Omega}}^{*}P^{s_{1}}\otimes{\bm{\Omega}}P^{s^{\prime}_{1}})({\bm{\Omega}}P^{s^{\prime}_{2}}\otimes{\bm{\Omega}}^{*}P^{s_{2}})({\bm{\Omega}}^{*}P^{s_{3}}\otimes{\bm{\Omega}}P^{s^{\prime}_{3}})\cdots({\bm{\Omega}}P^{s^{\prime}_{L}}\otimes{\bm{\Omega}}^{*}P^{s_{L}})].$ (116)
The corresponding local MPO representation of $K_{c}$ can then be derived as demonstrated in Sec. IV. Introducing the ${\mathbb{Z}}_{N}$ swap operator
$S=\frac{1}{N}\sum_{a,b=1}^{N}(X^{a}Z^{b})\otimes(X^{a}Z^{b})^{\dagger},$ (117)
we can rewrite Eq. (116) as
$\langle{\bf s}^{\prime}|K_{c}|{\bf s}\rangle$ $={\rm Tr}[({\bm{\Omega}}^{*}P^{s_{1}}\otimes{\bm{\Omega}}P^{s^{\prime}_{1}})S({\bm{\Omega}}^{*}P^{s_{2}}\otimes{\bm{\Omega}}P^{s^{\prime}_{2}})S\cdots({\bm{\Omega}}^{*}P^{s_{L}}\otimes{\bm{\Omega}}P^{s^{\prime}_{L}})S]$ (118)
which is represented by the site-independent local MPO tensor
$({\cal K}^{j}_{c})_{\alpha_{j-1},\alpha_{j}}$ $=\sum_{s_{j},s^{\prime}_{j}}[({\bm{\Omega}}^{*}P^{s_{1}}\otimes{\bm{\Omega}}P^{s^{\prime}_{1}})S]_{\alpha_{j-1},\alpha_{j}}|s^{\prime}_{j}\rangle\langle s_{j}|$
$=\omega^{h_{j-1}g_{j}-g_{j-1}h_{j}}|g_{j}\rangle\langle h_{j}|.$ (119)
Just as in Sec. IV, each bond index $\alpha_{j}=(g_{j},h_{j})$ is a pair of ${\mathbb{Z}}_{N}$ variables.
Recalling the NQS decomposition ${\bm{\Omega}}={\bf W}{\bf W}^{t},{\bm{\Omega}}^{*}={\bf W}^{*}{\bf W}^{\dagger}$, one can re-organize Eq. (115) as
$\langle{\bf s}^{\prime}|K_{c}|{\bf s}\rangle$ $={\rm Tr}[({\bf W}^{\dagger}P^{s_{1}}{\bf W})({\bf W}^{t}P^{s^{\prime}_{2}}{\bf W}^{*})({\bf W}^{\dagger}P^{s_{3}}{\bf W})\cdots({\bf W}^{t}P^{s^{\prime}_{L}}{\bf W}^{*})]\times{\rm Tr}[({\bf W}^{t}P^{s^{\prime}_{1}}{\bf W}^{*})({\bf W}^{\dagger}P^{s_{2}}{\bf W})({\bf W}^{t}P^{s^{\prime}_{3}}{\bf W}^{*})\cdots({\bf W}^{\dagger}P^{s_{L}}{\bf W})]$
$={\rm Tr}[({\bf W}^{\dagger}P^{s_{1}}{\bf W}\otimes{\bf W}^{t}P^{s^{\prime}_{1}}{\bf W}^{*})({\bf W}^{t}P^{s^{\prime}_{2}}{\bf W}^{*}\otimes{\bf W}^{\dagger}P^{s_{2}}{\bf W})\cdots({\bf W}^{t}P^{s^{\prime}_{L}}{\bf W}^{*}\otimes{\bf W}^{\dagger}P^{s_{L}}{\bf W})]$
$={\rm Tr}[({\bf W}^{\dagger}\otimes{\bf W}^{t})(P^{s_{1}}\otimes P^{s^{\prime}_{1}})({\bf W}\otimes{\bf W}^{*})\cdots({\bf W}^{t}\otimes{\bf W}^{\dagger})(P^{s^{\prime}_{L}}\otimes P^{s_{L}})({\bf W}^{*}\otimes{\bf W})]$
$={\rm Tr}[({\bf W}^{\dagger}\otimes{\bf W}^{t})(P^{s_{1}}\otimes P^{s^{\prime}_{1}})({\bf W}\otimes{\bf W}^{*})S\cdots S({\bf W}^{\dagger}\otimes{\bf W}^{t})(P^{s_{L}}\otimes P^{s^{\prime}_{L}})({\bf W}\otimes{\bf W}^{*})S].$ (120)
This gives rise to another MPO representation of $K_{c}$ with the local tensor
$({\cal K}^{j}_{c})_{\alpha_{j-1},\alpha_{j}}$ $=\sum_{s_{j},s^{\prime}_{j}}[({\bf W}^{\dagger}\otimes{\bf W}^{t})(P^{s_{j}}\otimes P^{s^{\prime}_{j}})({\bf W}\otimes{\bf W}^{*})S]_{\alpha_{j-1},\alpha_{j}}|s^{\prime}_{j}\rangle\langle s_{j}|$
$=\sum_{s_{j},s^{\prime}_{j}}W^{*}(s_{j},g_{j-1})W(s^{\prime}_{j},h_{j-1})W(s_{j},h_{j})W^{*}(s^{\prime}_{j},g_{j})|s^{\prime}_{j}\rangle\langle s_{j}|.$ (121)
A substantially different approach to the MPO representation of $K_{c}$ can be made by starting from the MPO representation of the ${\mathbb{Z}}_{N}$ KW operator in Eq. (63). Invoking the definition of the KW operator for the cluster state, $K_{c}=TK_{1}K_{2}$, we get the MPO for $K_{c}$ as
$K_{c}=T\sum_{\bf g}\bigotimes_{j}({\cal K}^{j})_{g_{j},g_{j+2}}=\sum_{\bf g}\bigotimes_{j}({\cal K}^{j}_{c})_{g_{j-1},g_{j+1}},$ (122)
where
$({\cal K}^{j}_{c})_{g_{j-1},g_{j+1}}$ $=T({\cal K}^{j-1})_{g_{j-1},g_{j+1}}$
$=|\overline{g_{j+1}-g_{j-1}}\rangle_{j}\langle g_{j-1}|_{j-1}.$ (123)
Note that each ${\cal K}^{j}_{c}$ contains $N\times N$ elements, but each element is an operator defined over the two adjacent qudits at $j$ and j1j-1. As a result, we have $K_{c}$ written as a product of two traces:
$K_{c}$ $={\rm Tr}[{\cal K}^{1}_{c}\otimes{\cal K}^{3}_{c}\otimes\cdots\otimes{\cal K}^{L-1}_{c}]$
$\times{\rm Tr}[{\cal K}^{2}_{c}\otimes{\cal K}^{4}_{c}\otimes\cdots\otimes{\cal K}^{L}_{c}].$ (124)
This can be re-written as a single trace
$K_{c}={\rm Tr}[{\cal K}^{(1,2)}_{c}\otimes{\cal K}^{(3,4)}_{c}\cdots\otimes{\cal K}^{(L-1,L)}_{c}$],
where the two adjacent sites are now grouped as one effective site and
$({\cal K}^{(2n-1,2n)}_{c})_{(g_{2n-2},g_{2n-1}),(g_{2n},g_{2n+1})}$
$=({\cal K}^{2n-1}_{c})_{g_{2n-2},g_{2n}}\otimes({\cal K}^{2n}_{c})_{g_{2n-1},g_{2n+1}}.$ (125)
Finally, we present yet another way to arrive at the MPO representation of $K_{c}$ by examining how the MPS representation of the cluster state wavefunction transforms under it:
$\Psi({\bf s})\xrightarrow{K_{c}}{\rm Tr}\left[\prod_{j=1}^{L}\left(\sum_{g_{j}}\omega^{g_{j}(s_{j+1}-s_{j-1})}A^{g_{j}}\right)\right]$ (126)
Performing the sum over each $g_{j}$ gives
$\sum_{g_{j}\in{\mathbb{Z}}_{N}}\omega^{g_{j}(s_{j+1}-s_{j-1})}A^{g_{j}}=\sum_{s^{\prime}_{j}}|s^{\prime}_{j}\rangle\langle\overline{s_{j+1}-s_{j-1}+(-1)^{j}s^{\prime}_{j}}|$ (127)
Note that the phase factor $\omega^{g_{j}(s_{j+1}-s_{j-1})}$ in this transformation can be understood as
$\omega^{g_{j}(s_{j+1}-s_{j-1})}=\langle s_{j-1}|({\cal K}^{j}_{c})_{s_{j-1},s_{j+1}}|g_{j}\rangle$ (128)
where the operator $({\cal K}^{j}_{c})_{s_{j-1},s_{j+1}}$ is defined as
$({\cal K}^{j}_{c})_{s_{j-1},s_{j+1}}$ $=\sum_{g_{j}}\omega^{(s_{j+1}-s_{j-1})g_{j}}|s_{j-1}\rangle_{j-1}\langle g_{j}|_{j}$
$=|s_{j-1}\rangle_{j-1}\langle\overline{s_{j+1}-s_{j-1}}|_{j}.$ (129)
It turns out that this gives rise to another MPO representation of $K_{c}$ that resembles Eq. (123):
$K_{c}=\sum_{\bf s}\bigotimes_{j}({\cal K}^{j}_{c})_{s_{j-1},s_{j+1}}.$ (130)
Appendix F Non-invertible symmetry of dCSII
The type-II dipolar cluster model admits the NIS [19]
$K_{d}=$ $\Big(\prod_{j}S_{2j,2j+1}\Big)\tilde{K}_{1}\tilde{K}_{2}^{\dagger}$
$=$ $\sum_{\mathbf{s},\mathbf{s}^{\prime}}\Phi(\mathbf{s},\mathbf{s}^{\prime})\Phi^{*}(\mathbf{s}^{\prime},\mathbf{s})|\mathbf{s}^{\prime}\rangle\langle\mathbf{s}|,$
$\Phi(\mathbf{s},\mathbf{s}^{\prime})=$ $\omega^{\sum_{j}s_{2j+1}(s^{\prime}_{2j+2}-2s^{\prime}_{2j}+s^{\prime}_{2j-2})}$
$=$ $\omega^{\sum_{j}(s_{2j-1}-2s_{2j+1}+s_{2j+3})s^{\prime}_{2j}}.$ (131)
Here, $\tilde{K}_{1}$ and $\tilde{K}_{2}$ are the dipolar KW operators or the quadrupole Fourier transformations
$\tilde{K}=\sum_{\mathbf{s},\mathbf{s}^{\prime}}\omega^{\sum_{j}(s_{j+1}-2s_{j}+s_{j-1})s^{\prime}_{j}}|\mathbf{s}^{\prime}\rangle\langle\mathbf{s}|$ (132)
acting on the odd and even sublattices, respectively. The swap operator $S_{2j,2j+1}$, already introduced in (117), interchanges the qudit states at $2j$ and $2j+1$.
The $K_{d}$-invariance of the dCSII ground state can be shown as follows:
$\Psi(\mathbf{s})\xrightarrow{K_{d}}$ $\sum_{\bf g}\Phi(\mathbf{g},\mathbf{s})\Phi^{*}(\mathbf{s},\mathbf{g})\Psi(\mathbf{g})$
$=\sum_{\mathbf{g}_{1}}\omega^{\sum_{j}g_{2j+1}(s_{2j+2}-2s_{2j}+s_{2j-2})}$
$\mkern 21.0mu\times\sum_{\mathbf{g}_{2}}\omega^{\sum_{j}g_{2j}(g_{2j-1}-2g_{2j+1}+g_{2j+3}-s_{2j-1}+2s_{2j+1}-s_{2j+3})}$
$\propto\sum_{k\in{\mathbb{Z}}_{N}}\omega^{\sum_{j}(s_{2j+1}+k)(s_{2j+2}-2s_{2j}+s_{2j-2})}$
$\propto\Psi(\mathbf{s}).$ (133)
In the third line, we used that the summation over the even sites $\mathbf{g}_{2}=(g_{2},g_{4},\dots,g_{L})$ gives the constraint
$g_{2j-1}-2g_{2j+1}+g_{2j+3}\equiv s_{2j-1}-2s_{2j+1}+s_{2j+3}~({\rm mod}~N)$
$\Longleftrightarrow g_{2j-1}=s_{2j-1}+k,\,k\in{\mathbb{Z}}_{N}.$
The definition of $\Phi({\bf s},{\bf s}^{\prime})$ makes clear that $K_{d}$ is a quadrupole Fourier transformation whose image lies entirely in the sector $C_{1}=C_{2}=D_{1}=D_{2}=1$.
Since $\Phi({\bf s},{\bf s}^{\prime})=\Psi(s_{1},s^{\prime}_{2},s_{3},\dots,s^{\prime}_{L})$, each matrix element of $K_{d}$ is nothing but the product of the dCSII wavefunction and its complex conjugate,
$\langle{\bf s}^{\prime}|K_{d}|{\bf s}\rangle=\Psi(s_{1},s^{\prime}_{2},s_{3},\dots,s^{\prime}_{L})\Psi^{*}(s^{\prime}_{1},s_{2},s^{\prime}_{3},\dots,s_{L}).$ (134)
This shows that the $K_{d}$-invariance of the dCSII wavefunction immediately implies the first four of the following fusion rules satisfied by the symmetry generators $\{K_{d},C_{1},C_{2},D_{1},D_{2}\}$ [19]:
$C_{1}K_{d}=K_{d}C_{1}=K_{d},\quad C_{2}K_{d}=K_{d}C_{2}=K_{d},$
$D_{1}K_{d}=K_{d}D_{1}=K_{d},\quad D_{2}K_{d}=K_{d}D_{2}=K_{d},$
$K_{d}^{\dagger}K_{d}=K_{d}^{2}=\left(\sum_{m=1}^{N}C_{1}^{m}\right)\left(\sum_{n=1}^{N}C_{2}^{n}\right)\left(\sum_{p=1}^{N}D_{1}^{p}\right)\left(\sum_{q=1}^{N}D_{2}^{q}\right).$ (135)
Using the TPS representations of the dCSII wavefunction (44) and (III.2), it is possible to write an MPO representation of $K_{d}$,
$K_{d}={\rm Tr}[{\cal K}_{d}^{(L,1)}\otimes{\cal K}_{d}^{(2,3)}\otimes\cdots\otimes{\cal K}_{d}^{(L-2,L-1)}],$ (136)
where the local tensor over each pair of adjacent sites $(2n,2n+1)$ is defined as
$\big({\cal K}_{d}^{(2n,2n+1)}\big)_{\Gamma_{n-1},\Gamma_{n}}=\sum_{\begin{subarray}{c}s_{2n},s_{2n+1}\\ s^{\prime}_{2n},s^{\prime}_{2n+1}\end{subarray}}\sum_{\beta,\beta^{\prime}}T^{s^{\prime}_{2n}}_{\alpha_{n-1}\beta\gamma_{n}}T^{s_{2n+1}}_{\alpha_{n}\beta\gamma_{n-1}}\Big(T^{s_{2n}}_{\alpha^{\prime}_{n-1}\beta^{\prime}\gamma^{\prime}_{n}}\Big)^{*}\Big(T^{s^{\prime}_{2n+1}}_{\alpha^{\prime}_{n}\beta^{\prime}\gamma^{\prime}_{n-1}}\Big)^{*}|s^{\prime}_{2n},s^{\prime}_{2n+1}\rangle\langle s_{2n},s_{2n+1}|,$ (137)
and each bond index $\Gamma_{j}=(\alpha_{j},\gamma_{j},\alpha^{\prime}_{j},\gamma^{\prime}_{j})$ is a collection of four ${\mathbb{Z}}_{N}$ numbers.